### v3

# Practical privacy metrics for synthetic data

Abstract
This paper explains how synthpop package111see www.synthpop.org.uk  [^2] The new version 1.9-0 is currently available on Github at https://github.com/gillian-raab/synthpop has been extended to include functions to calculate identity and attribute disclosure risk measures for the original records. using information obtained from the corresponding synthetic data. The basic function disclosure calculates identity disclosure for a set of quasi-identifiers (keys) and attribute disclosure for one variable specified as a target from the same set of keys. The second function multi.disclosure is a wrapper for the first and it presents summary results for a set of targets. This short paper explains the measures of disclosure risk and documents how they are calculated. We recommend two measures: $RepU$ (replicated uniques) for identity disclosure and $DiSCO$ (disclosive in synthetic and correct in the original) for attribute disclosure. Both are expressed a % of the original records and both can be compared to similar measures calculated from the original data. Experience with using the functions on real data found that some apparent disclosures could be identified as coming from relationships in the data that would be expected to be known to anyone familiar with its features. We flag cases when this seems to have occurred and provide means of excluding them.
1 Introduction
In his recent review of thirty years of synthetic data Reiter [17] comments:
”While there is need to examine disclosure risks in synthetic data, there is no standard for doing so, especially in fully synthetic data. Instead, disclosure checks tend to be ad hoc”
This is in contrast to the variety of measures of utility available for synthetic data; see [16] for a list of these. While utility measures must be chosen that are relevant to the intended use of the synthetic data, disclosure measures must focus on the possible harm to the privacy of an individual or other unit whose data contributed to the creation of the synthetic data. The evaluation of the disclosure risk from synthetic data must relate to the context of its release (see [5] for a discussion of this). We cannot expect a fixed rule, for example that a criterion for release requires a value of some disclosure measure beyond a threshold. Instead, we expect those releasing data to use the disclosure measures to evaluate potential harm to the data subjects or to the data custodians from information in the released data. We hope that these new functions will allow the disclosure risk of synthetic data sets to be explored and, where necessary, reduced.
We expect that disclosure risks from synthetic data to be lower than those from the original data. But in some cases, e.g. a large data set with only a few categorical variables, the disclosure risk from the original may already be low. To evaluate the disclosure risk of a synthesis method, the risks for synthetic data must be compared with equivalent risks for the original. We consider two types of disclosure risk:
  • identity disclosure: This refers to the ability to identify individuals in the data from a set of known characteristics that we will refer to as keys. Identity disclosure may be less relevant for completely synthesized [^3] Complete or full synthesis is when all values of all variables are replaced by synthetic values. This is in contrast to incomplete or partial synthesis where only some variables are replaced. data because there is no one-to-one correspondence between records in the original and synthetic data sets. But it may still be of interest since it is an important factor for attribute disclosure.
  • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the original data.
The measures we describe here are appropriate for fully synthetic data where all items in all records are replaced by synthetic values. Different measures have been developed for partially synthetic data [3, 18]
The disclosure risk posed by synthetic data can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques. These methods can be used to reduce disclosure risk by modifying the original data before synthesis, or the synthetic data before its is released. Some such methods are already available in the synthpop package. These include categorising, top/bottom coding and smoothing for continuous variables, and the merging of small categories for factors. The removal of replicated uniques is another option available.
There has been many recent proposals for making synthetic data sets Differential Private (DP) (selection of refs). DP is a very strong privacy guarantee that protects against an intruder with arbitrary external information about the subjects in the data, except for the one whose privacy is being protected. This is an unrealistic assumption and DP synthetic data has been shown to have unacceptably low utility in many cases [1, 20]. We will not discuss these methods here, but note that we could use the metrics proposed here to evaluate disclosure risks for DP synthetic data.
2 A simple example
Here we illustrate the basic use of multi.disclosure. If the parameter targets is not specified, all the variables in the synthetic data that are not part of keys are used as targets. The identity disclosure measures are UiO for original and repU for synthetic, and for attribute disclosure Dorig for original and DiSCO for synthetic. These and other measures will be explained in Sections 3.2 and 3.3.
First, a subset of 9 variables are selected fron the SD2011 data (a survey on quality of life in Poland) that is available as part of the synthpop package. A single synthetic data set is created by the default method in synthpop: cart for each conditional distribution. The synthetic data object s1444an object of class synds has a component syn that is a single synthetic data set. The disclosure functions can also be used with synthetic data created by other methods either as single synthetic data sets or lists of repeated syntheses from the same original. Here we select 4 keys that represent items that might be known for members of this sample, or of the Polish population in 2011. By default,the 5 other variables become the targets: income (rounded to give 406 distinct values) marital (marital status), workab (intention to work abroad), ls and depress (scores for life satisfaction and depression with 7 and 21 distinct values). We first use multi.disclosure to create a multi.disclosure object for these keys, and we print out its identity component.
R> library("synthpop")
R> ods <- SD2011[, c("sex", "age", "region","placesize","depress",
+    "income","ls","marital" , "workab")]
R> s1 <- syn(ods, seed = 8564, print.flag = FALSE)
R> t1 <- multi.disclosure(s1, ods, print.flag = FALSE, plot = FALSE,
+     keys = c("sex", "age", "region", "placesize"))
R> print(t1,to.print = "ident")
Disclosure risk for 5000 records in the original data
Identity disclosure measures
from keys: sex age region placesize
For original  ( UiO )  48.38 %
For synthetic ( repU ) 14.86 %.
The measure $UiO$ (Unique in Original) shows that 48% of the original records would have unique combinations of these 4 keys. The term, ”singling out” is used in data protection regulation for this type of attribute disclosure [^5] See for example its use in the UK Information Commissioner’s guidance on anonymisation here https://ico.org.uk/media/about-the-ico/documents/4018606/chapter-2-anonymisation-draft.pdf. For the synthetic data $RepU$ tells us that almost 15% of the original records would be unique in the original and also unique in the synthetic data.
For attribute disclosure we can examine the results as either a table or a plot. Here we see in Figure 1 the plot that would have been generated if plot had been set to TRUE in the code above, the attribute disclosure results would have been plotted as shown in Figure 1. Details of how identity and disclosure measures are calculated can be found in Section 3.
Figure 1: Plot from multi.disclosure with plot = TRUE, the default value.
The results in the plot are in descending order by the disclosure in the synthetic data. The measure $Dorig$ tells us that these 4 keys would identify a unique value of each of these targets for over 50% of the original records. $DiSCO$, the proportion of the original records that are disclosive in the original and also in the synthetic data with a correct attribution to the target give lower values. The first two variables shown in Figure 1 have additional labels that flag possible contributions to disclosure from knowledge of 1-way or 2-way relationships in the original data. This is detailed in Section 4.
Previous work on attribute disclosure [4, 22] has used the Correct Attribution Probability ($DCAP$) as a disclosure measure for synthetic data. This is calculated as the percentage of records correctly predicted from the synethetic data by a set of keys. However it has been suggested [7, 2], that it may be as much a measure of utility as disclosure risk and modifications have been suggested [21, 10] one of which ($TCAP$) is close to our proposed measure, $DiSCO$. For completeness the measures $DCAP$ and $TCAP$ are calculated by disclosure; see Appendix 2.
3 Scenario and definitions
3.1 Setting the scene
These disclosure measures are intended to assess what a person who only has access to the synthetic data can infer about known individuals who are present in the original data. We will use the term ”intruder” for such a person, though no malicious intent is implied. The intruder is assumed to have information for one or more individuals about the value of certain key variables that are present in the same format in the original and the synthetic data. They attempt first to see if the individual is present, and then to determine the value of other items in the data file that we refer to as targets. We are assuming a worst-case scenario where the intruder believes they are querying the original data. [^6] This may not be too unrealistic if the data are made available inadvertently, or if the intruder thinks that efforts to label the synthetic data as e.g. ”Fake Data” are thought just to be a cover up. Disclosure measures from the synthetic data are each compared to similar measures for someone with access to the original data. Here we will introduce the measures by an example. Formal definitions with notation and formulae are in Appendix 1. The first step in evaluating disclosure risk, as described here, is to identify a set of keys that might be expected to be known to an intruder. These keys are then combined to form a quasi-identifier that we designate as $q$. For example, if we have hospital records we might define age,sex date and hospital as keys and this would give a $q$ with levels such as ”78 | M | 1/1/2024 | WG” for a 78 year old man admitted to hospital WG on 1/1/2024.
3.2 Identity disclosure measures
The concept of k-anonymity is central to identity disclosure for microdata. First proposed in 1998 [19] it is discussed fully in [5]. A table is k-anonymous if a set of keys identifies at most $k-1$ individuals. Thus 2-anonymous data will never identify just one individual. Based on this idea, the percentage of records for which the keys identify just one individual give identity disclosure measures. A table of $q$ values is produced from the synthetic and the original data. $UiO$ and $UiS$ are the percentages of records with keys where the table value is 1. An intruder checking out a record for their known set of keys will look for it in the synthetic data. Some records will not be in the synthetic data and $UiOiS$ (Unique in Original in Synthetic) gives the % that would be found. These records are then checked for uniqueness in the synethetic data giving, $repU$ is the percentage of unique original records that are also unique in the synthetic data.
The percentage $repU$ has been used as a disclosure measure to evaluate synthetic data by [9] and by [15]  [^7] Jackson et al. in [9] argue that the denominator for $repU$ should be $N_{s}$ rather than $N_{d}$. This is inappropriate because our scenario is to consider the risk to the original data.. Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the synthetic data. Nowok et al. [13] have evaluated this and give an example where this process has very little effect on utility. The function replicated.uniques [^8] For example by replicated.uniques (s2, ods, keys = c("sex","region","age","placesize") also calculates $repU$ using a different method from the one described here.
One of the outputs of the functions disclosure and multi.disclosure is ident, a table of identity disclosure measures as illustrated by this example, using the keys from the example in Section 2 but now calculated for a synthetic object with 5 data sets and using the one target depress. We first create t5 an object of class disclosure and then print out the identity and attribute disclosure measures for each synthetic data set.
The table for identity disclosure has $UiO$ and $repU$ as its first and last column. The 2nd and 3rd columns are $UiS$ calculated from the synthetic data in the same manner as $UiO$ from the original and $UiOiS$ the % of $UiO$ with $q$ that are in the synthetic data but not necessarily unique. Each are steps towards calculating $repU$.
R> s5 <- syn(ods, seed = 8564, m = 5, print.flag = FALSE)
R> t5 <- disclosure( s5, ods, keys = c("sex", "age", "region",
+    "placesize"), target = "depress", print.flag = FALSE)
R> print(t5, to.print = c("ident"))
Disclosure measures from synthesis for 5000 records in original data.
Identity disclosure measures for 5 synthetic data set(s) from keys:
 sex age region placesize
    UiO   UiS UiOiS  repU
1 48.38 37.34 22.68 14.86
2 48.38 35.44 22.24 13.96
3 48.38 35.18 21.98 13.62
4 48.38 34.90 22.08 13.78
5 48.38 36.14 22.00 14.62
3.3 Attribute disclosure measures
We approach disclosure risk from the point of view of an intruder with access to the synthetic data and to certain attributes (quasi-identifiers) known for one or more individuals in the original data. The quasi-identifiers for the known record(s) can be combined to create a composite variable $q$ that can be created for all records in the original and synthetic data. Modelling what an intruder might do we calculate the following measures, each of whuch is a proportion of the original records:
  • Look up $q$ for each original record in the synthetic data. The proportion found becomes $iS$ (in Synthetic).
  • Check if all records with the same $q$ have the same level of the target $t$. The proportion passing this further test becomes $DiS$ (Disclosive in Synthetic).
  • Then check if these apparent disclosures corresponds to the value of $t$ in the original data. The proportion of original records for which this is true becomes $DiSCO$ Disclosive in Synthetic Correct in Original.
Note that records contributing to $DiSCO$ may not be disclosive in the original data, as this information would not be available to the intruder. A further measure $DiSDiO$ (Disclosive in Synthetic Disclosive in Original) restricts the score to those disclosive in the original. These measures are defined formally in Appendix 1.
R> print(t5, to.print = c("attrib"))
Disclosure measures from synthesis for 5000 records in original data.
Attribute disclosure measures for depress from keys: sex age region placesize
  Dorig  Dsyn    iS   DiS DiSCO DiSDiO max_denom mean_denom
1  53.3 46.26 64.90 34.18  9.54   6.14         3       1.16
2  53.3 44.80 64.00 32.50 10.26   6.78         4       1.19
3  53.3 44.60 64.02 32.14  9.10   5.92         4       1.19
4  53.3 45.80 63.88 33.38  9.20   5.52         4       1.21
5  53.3 44.52 63.44 31.46  9.34   5.80         4       1.23
Here we print the table of attribute disclosure measures for the example in the previous section. As we move from $iS$ to $DiS$ and to $DiSCO$ we can see how the %disclosive is affected by different conditions. Here lack of $q$ levels in the synthetic data retains just 64% of records. The requirement for a record to be disclosive in the synthetic data reduces this to 34% and again to around 9% for those with a corect attribution. This reduces it to around 6% by restricting to records disclosive in the original.
The $Dorig$ and $DiSCO$ measures are not restricted to disclosures that are identified from unique records for $q$ in either the original or the synthetic data. The number of records contributing to each disclosive $qt$ cell in the original table is the denominator that applies to that record. The columns max denom and mean denom for the records disclosive in the synthetic data contributing to the $DiSCO$ refer to the denominators ($d_{.q}$ used to calculate $pd_{tq}$ from the original data). We can see from the mean that here the majority of disclosive records had unique key combinations in the original data, and the maxima was 3 for the first synthesis and 4 for the others. Large denominators can be an indication that some of the disclosure may be coming from strong relationships between variables in the data that might even be expected a-priori. This aspect is discussed further in Section 4. The disclosure measures can be restricted to those with small denominators by using the parameter exclude_over_denom_lim to TRUE. The example in Section 2 is here run restricted to denominators of 1.
R> multi.disclosure(s1, ods, print.flag = FALSE, plot = FALSE,
+    keys = c("sex", "age", "region", "placesize"),
+    denom_lim =1, exclude_ov_denom_lim = TRUE)
Disclosure risk for 5000 records in the original data
Identity disclosure measures
from keys: sex age region placesize
For original  ( UiO )  50.26 %
For synthetic ( repU ) 16.36 %.
Table of attribute disclosure measures for sex age region placesize
Original measure is  Dorig and synthetic measure is DiSCO
Variables Ordered by synthetic disclosure measure
                     attrib.orig attrib.syn check1 Npairs check2
1 income                   48.38       4.02             0
2 depress                  48.38       7.02             0
3 ls                       48.38       9.74             0
4 marital                  48.38      15.72             0
5 workab 1way checks       48.38      20.62     NO      0
The $DiSCO$ values have decreased, as expected, but $UiO$ and $repU$ have increased and also that $Dorig$ is now the same for all targets and equal to the $UiO$ value before the denominator exclusion. This makes sense because removing large denominators from $UiO$ and $repU$ increases the number of uniques. Also, with denominators of 1, all unique values of $q$ are disclosive in the original data.
Note that decreases in $DiSCO$ are more pronounced for workab and marital, the two targets that were flagged to check 1-way or 2-way relationships. Their original $DiSCO$ values were 37% for marital and 53% for workab. This and other methods of excluding certain apparent disclosures are discussed in the next section.
4 Identifying disclosure from 1-way and 2-way relationships
As mentioned in the Introduction, what we can learn about disclosiveness of attributes can depend on our prior knowledge of the data set or the population from which it is drawn. It would not be practical to specify our prior probability for every possible combination of keys. However, checking two aspects of disclosure results can help us to check when we might have predicted the correct attribution with high probability without knowledge of all the keys. The first is a check for a target where a high proportion of records have one level of the target. The second is when there is a strong relationship between the target and one of the keys, so that one $tq$ pair accounts for many of the disclosive records. These two aspects are flagged by the values of check_1way and check_2way that are returned as part of objects of Returning to the example in 2 we now use the function disclosure to get details for the two targets that were flagged as requiring checking; see Figure 1.
R> d1_workab <- disclosure(s1, ods, print.flag = FALSE, target = "workab",
+     keys = c("sex", "age", "region", "placesize"),plot = FALSE)
R> print(d1_workab, to.print = c("check_1way","check_2way"))
Disclosure measures from synthesis for 5000 records in original data.
Details of target level contributing disproportionately to disclosure
  check Level totalDisclosive PctDisAll totalLevel pctdisLevel
1     1    NO            2661     53.22       2529       95.04
Details of target-key combinations contributing disproportionately to disclosure
This is a list with one for each of syntheses
[[1]]
  total pctall max_denom    key        target_key pct_2way denom_2way
1     6    0.2         6    age           NO | 55     97.1        105
2     6    0.2         6 region NO | Podkarpackie     95.2        313
We can see that it was the category ”NO” of workab that contributed most to the disclosure risk; most survey respondents (89%) had never worked abroad. This level of the target accounted for 93% of apparent disclosures for workab. To predict this level for all of a group with the same $q$ could hardly be considered disclosive. This would be true if the intruder had access to the marginal distribution of workab, and even if they did not some knowledge of the respondents’ background might suffice. The $tq$ pairs identified for workab both included the ”NO” level of the target.
[…]
R> d1_marital <- disclosure(s1, ods, print.flag = FALSE, target = "marital",
+     keys = c("sex", "age", "region", "placesize"),plot = FALSE)
R> print(d1_marital, to.print = c("check_1way","check_2way"))
Disclosure measures from synthesis for 5000 records in original data.
No 1 way checks for your target are flagged at your settings for thresh_1way
Details of target-key combinations contributing disproportionately to disclosure
This is a list with one for each of syntheses
[[1]]
  total pctall max_denom key   target_key pct_2way denom_2way
1     7    0.4         7 age MARRIED | 40     82.3         79
2     6    0.3         6 age  SINGLE | 20    100.0         79
For the target marital two $tq$ pairs are identified as contributing large denominators. Those aged 20 are all SINGLE, and those aged 40 are mainly MARRIED. Again these would not be considered disclosive for an intruder with some knowledge of the respondents.
The thresholds for identifying one-way and two-way relationships can be modified. In each case there are two criteria that need to be satisfied, one number and one %. For one-way disclosure the parameter thresh_1way has the default value of $c(50,90)$, meaning that there must be at least 50 disclosive records for one level of a target and that $q$ values including this target must account for over 90% of all disclosive records. For two way relationships the thresh_2way has default value $c(5,80)$ and the algorithm first identifies all disclosive $tq$ combinations with denominators over 5. It then identifies the level of the key in $q$ with that best predicts this level of $t$ and checks if over 80% of the disclosive records with this key would have the correct prediction.
5 Excluding records
Having identified records where the disclosure is not really telling us anything new, one option is to exclude some of these original records explicitly from the measures. The following parameters for multi.disclosure and disclosure can be used for this.
  • not.target: All records with levels of the target given by the parameter not.target are excluded from all disclosure measures.
  • usekeysNA:This is set to TRUE by default so that missing values are included in all tables. It can be set to FALSE for some or all keys to exclude NAs.
  • usetargetsNA)(multi.disclosure) or usetargetNA (disclosure) ): Similar to the above for target(s) - note can be a vector for multi.disclosure.
  • exclude.keys, exclude.keylevs and exclude.targetlevFor disclosure only: Three vectors of length the number of key-target pairs to be excluded from all tables.
Figure 2: Plot from multi.disclosure for variables flagged by checks.
To illustrate exclusions we will use the Adult data set from the UCI machine learning repository [12] with almost 50 thousand records from the US Census income study, available as part of the arules package for R [8]. This is one of the data sets used in [6] to evaluate privacy risks for synthetic data. Figure 2 is the output of multi.disclosure from the keys age, occupation, race and sex for the other 10 variables in the data. The code used to calculate the results in this section is in Appendix 3. The disclosiveness of the original data is relatievly low, compared to the previous example, because of the large sample size and the absence of any geographic identifiers. Three of the 10 variables capital.gain, capital.loss and native.country are flagged to check one-way relationships. The check1 column of the attribute table tells us that the levels contributing to disclosre are zeros for the first two and United-States for native.country. These levels make up 95%, 92% and 89% of all records.
Three other variables are flagged as having disclosive two_way relationships workclass, marital, relationship with totals of 7,7 and 2 $tq$ combinations respectively. The summary function multi.disclosure does not allow target and key specific pairs to be excluded. To investigate this it is necessary to examine the output of disclosure999see code in Appendix 3 for how to do this.. This showed that the largest contribution to check_2way for workclass was due to this being missing when occupation was missing, although other relationships between these two variables also contributed.
denom_lim 1
NAs out NAs out
No excl not.tlev not.tlev not.tlev denom_lim 1
target orig syn orig syn orig syn orig syn orig syn
capital.gain 22.55 19.87 0.21 0.00 0.21 0.00 0.19 0.00 2.68 1.17
capital.loss 30.61 27.71 0.08 0.00 0.08 0.00 0.08 0.00 2.68 1.26
education 3.71 0.93 3.71 0.93 3.71 0.93 2.45 0.46 2.68 0.50
fnlwgt 2.70 0.00 2.70 0.00 2.70 0.00 2.45 0.00 2.68 0.00
hours.per.week 4.36 0.19 4.36 0.19 4.36 0.19 2.45 0.12 2.68 0.13
income 4.97 2.10 4.97 2.10 4.97 2.10 1.58 0.50 2.68 0.82
marital.status 8.23 5.27 8.23 5.27 8.23 5.27 2.45 0.68 2.68 0.74
native.country 17.09 13.55 0.94 0.08 0.94 0.08 0.67 0.04 2.68 0.87
relationship 5.17 2.64 5.17 2.64 5.17 2.64 2.45 0.69 2.68 0.73
workclass 14.27 11.64 14.27 11.64 14.27 11.64 2.45 0.85 2.68 0.96
Table 1: Disclosure results from Adult data with different exclusions.
Table 1 gives the results of excluding different entries in the tables of $q$ and $t$ from the attrinute disclosure measures. Excluding the levels of the target flagged by textttcheck_1way reduces the $DiSCO$ to almost zero for these 3 variables. Adding exclusion of missing values reduced the disclosure for workclass and for some other variables a little. Adding the restriction to denominators of 1, reduced the disclosure for variables identified by check_2way to low levels. In the final columns we can see that restricting to denominators of 1, by itself, gives low levels of disclosure for all variables, with the exception of those flagged by check_1way. This approach was one that was trialed in an earlier version of these functions but discarded for giving misleading results for some data. It does not count some attribute disclosures that might well be found by an intruder. For example, if a small number of records with the same keys all had the same level of the target that corresponded to the level in the original data. We feel that a better approach is to exclude specific target/key combinations.
6 Conclusions
The privacy metrics we propose here are in some sense the opposite of differential privacy (DP). DP claims to protect data from an intruder with arbitrary knowledge of the data, except of the one record that has the greatest influence on the likelihood of the results. In contrast our metrics require the user to specify the details of what they would expect an intruder to know about the data. Most importantly, the user must specify keys that identify variables in the data that would expect to be known about individuals. Also, the routines flag cases where part of the disclosure measures come from one- or two-way relationships so that disclosure would be expected even in the absence of data. We have evaluated the routines on a few data sets and set levels of the thresholds for this at what seem to be reasonable levels, but more experience with other data sets would be valuable.
We hope that these tools will be helpful to data holders who need to make decisions about the risks of releasing synthetic data to the public, or to a restricted audience. They should also enable the disclosure risks of different synthesis methods to be evaluated.
7 Acknowledgement
Research Data Scotland (https://www.researchdata.scot/ funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package101010 from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control. We also thank the Scottish Centre for Administrative Data Research for continue to support the development of the synthpop package since its creation was supported by the ESRC funded SYLLS poject in 2012-14. We would also like to thank Timon Huijser for noting some mistakes in an earlier version of this paper that are now corrected
Appendix 1: Notation and formal definitions
Before defining the measures of identity and disclosure risk we need to introduce the notation that will be used to calculate them. The first step is to create the quasi-identifiers from the keys for the original and synthetic data. For the keys used in the example given in Section 2 the quasi-identifier that we will designate as $q$ for the first record in the original data is:
"FEMALE | 57 | Lubuskie | URBAN 100,000-200,000"
and that for the first record in the synthetic data:
"FEMALE | 39 | Zachodnio-pomorskie | URBAN 100,000-200,000".
In order to calculate identity disclosure measures, we need to compare the tables of $q$ from the original and synthetic data. For attribute disclosure measures we need to cross-tabulate $q$ with each target variable $t$ and compare findings from the synthetic data with what would have been found from the original data. In general, the levels of $q$ and sometimes $t$ in the original and synthetic data will not be the same. Before creating any tables, we need to define sets of $q$ and $t$ values that give the union of both sets of levels and align the tables so that their indices correspond.
For the original data $d_{.q}$ is the count of records with the keys corresponding to the levels of $q$ and $q_{tq}$ the count of records with this $q$ and level $t=1,...T$ of the target. The equivalent counts from the synthesised data are designated by $s_{.q}$ and $s_{tq}$. When a member of $q$ is in the original data but not in the synthetic, $s_{.q}$ and $s_{tq}$ are all zero. Similarly when a member of $q$ is in the synthetic data but not in the original, $d_{.q}$ and $d_{tq}$ are all zero. The two tables can be written as shown in Table 1, where the total records in the original data is $N_{d}$, made up of $N_{d~{}only}$ and $N_{d~{}both}$. The equivalent totals for the synthetic data are $N_{s}$, $N_{s~{}only}$ and $N_{s~{}both}$.
only in original in both only in synthetic Total
1 $d_{1q}$ $d_{1q}$ 0 $d_{1.}$
t $d_{tq}$ $d_{tq}$ 0 $d_{t.}$
T $d_{Tq}$ $d_{Tq}$ 0 $d_{T.}$
Column sums $d_{.q}$ $d_{.q}$ 0 $N_{d}$
Totals $N_{d\>only}$ $N_{d~{}both}$ 0 $N_{d}$
only in original in both only in synthetic Total
1 0 $s_{1}q$ $s_{1}q$ $s_{1.}$
t 0 $s_{t}q$ $s_{t}q$ $s_{t.}$
T 0 $s_{T}q$ $s_{T}q$ $s_{T.}$
Column sums 0 $s_{.q}$ $s_{.q}$ $N_{s}$
Totals 0 $N_{s~{}both}$ $N_{s\>only}$ $N_{s}$
Table 2: Notation for tables from quasi-identifier ($q$) and target ($t$) from original (upper table) and synthetic data (lower table).
To calculate the % of records in the original and synthetic data we need:
%$\%~{}Unique~{}in~{}Original=UiO=100\sum{(d_{.q}|d_{.q}=1})/N_{d}.$ (1)
%$\%~{}Unique~{}in~{}Synthetic=UiS=100\sum{(s_{.q}|d_{.q}=1})/N_{d}.$ (2)
The intruder has information about the keys for an individual in the real data that they attempt to identify in the synthetic data. They first attempt to find them in the synthetic data, abd the % found is:
%$\%~{}Unique~{}in~{}Original~{}in~{}Synthetic=UiOiS=100\sum{(d_{.q}=1|s_{.q}=1\land d_{.q}>0})/N_{d}.$ (3)
Some of these records would not be unique in the synthetic data, restricting to such records gives:
%$\%~{}replicated~{}Uniques=repU=100\sum{(s_{.q}|d_{.q}=1\land s_{.q}=1)}/N_{d}.$ (4)
To find an attribute from a set of keys, it is necessary to examine the distribution of $s_{tq}$ for groups defined by $q$. We define column proportions for the original and synthetic data as $pd_{tq}=d_{tq}/d_{.q}$ and for the synthetic as $ps_{tq}=s_{tq}/s_{.q}$.
Returning to the scenario described in Section 3.1, we must first define a measure of attribute disclosure for the original data. This is based on the concept of l-diversity [11] that requires that each set of records defined by $q$ has at least $l(\geq{2}$ distinct values of the target. A data set is l2-diverse for $q$ and $t$ if all records for every $q$ have the same level of $t$ [^11] This could be generalised to $l>2$, but in practice the levels of targets are not generally exchangeable and a more practical approach would be to aggregate levels for certain targets. An attribute disclosure measure for the original data can be defined as % Disclosive in Original :
$Dorig=100\sum^{q}{\sum^{t}{(d_{tq}|pd_{tq}=1})}/N_{d}.$ (5)
The equivalent measure for the synthetic data, taken as if it were real, becomes :
$Dsyn=100\sum^{q}{\sum^{t}{(s_{tq}|ps_{tq}=1})}/N_{s}.$ (6)
An intruder with access only to the synthetic data, but with knowledge of $q$ from one or more individuals in the original, would look them up in the synthetic data. Some of their $q$ levels be key combinations that do not appear in the synthetic data leaving the proportion that do appear as $iSO$ (in in Synthetic Original)
$iSO=100\sum^{q}{\sum^{t}{(d_{tq}|s_{tq}>0})}/N_{d}.$ (7)
A level of $q$ from an original record may identify more than one target in the synthetic data, or identify the wrong target. To exclude these we require the records to be Disclosive in the Synthetic data and Correct when checked with the original giving:
$DiSCO=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1})}/N_{d}.$ (8)
Note that $DiSCO$ can include records that are not disclosive in the original data giving a further measure Disclosive in Synthetic and Disclosive in the Original :
$DiSDiO=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1\land pd_{tq}=1})}/N_{d}.$ (9)
As we comment above the intruder would not be able to tell if records were identified as $DiSDiO$ rather than $DiSCO$, so we prefer the latter measure. However, the intruder can identify when the apparently disclosive record is not unique in the synthetic data. This restriction can be imposed by requiring the denominator in the synthetic data to exceed a 1, as described and discussed in Section 5.
Appendix 2: CAP measures
The measures $baseCAPd$, $DCAP$, $TCAP$ are calculated by the function disclosure and are stored in the component allCAPs of the output object of class disclosure. This is printed when the parameter to.print includes allCAPs. The first measure is known as the baseline CAP and refers to an average of the predictions that would be made by someone who only has access to the marginal distribution of the target. The intruder then guesses the CAP for each level of the target according to the relative frequencies $pd_{t.}$. Averaging this over all observations gives
$baseCAPd=\sum{(pd_{t.})^{2}}$
The % of $t$ that would be correctly predicted from $q$ for someone with access to the original data is $CAPd$:
$CAPd=\sum\limits_{tq}{(pd_{tq}}d_{tq})/N_{d},$
and the equivalent measure for the synthetic data treated as if it were the original is
$CAPd=\sum\limits_{tq}{(ps_{tq}}s_{tq})/N_{s}.$
The measure DCAP is the percentage of $t$ correctly predicted in the synthetic data $q$. This gives
$DCAP=\sum\limits_{tq}{(ps_{tq}}d_{tq})/N_{d}$
The $DCAP$ score can be expressed by scaling it from $baseCAPD$ to 1[10, 14], although this can result in some negative values. The $TCAP$ measure as defined by [10] uses the same count of disclosive records as $DISCO$, but uses as denominator the number of records that have keys represented in both the synthetic and the original data. It can also be scaled from $baseCAPD$ to 1, and as [10] demonstrate this measure can also take negative values.
R> print(t5, to.print = "allCAPs")
Disclosure measures from synthesis for 5000 records in original data.
CAP measures for the target depress with keys
sex age region placesize
  baseCAPd  CAPd  CAPs  DCAP  TCAP
1     9.81 74.15 69.78 16.39 14.70
2     9.81 74.15 69.36 17.45 16.03
3     9.81 74.15 69.24 16.20 14.21
4     9.81 74.15 69.87 15.92 14.40
5     9.81 74.15 69.00 16.17 14.72
As expected the lower denominator for $TCAP$ gives values higher than the equivalent $DiSCO$ values for the same example given in Section 3.3.
Appendix 3:: Code for analysis of the Adult data set.
library(arules)
data(AdultUCI)
names(AdultUCI) <- gsub("-",".", names(AdultUCI))
dim(AdultUCI)
##------ drop one of two identical education variables---------------------------
adultsample <- AdultUCI[,-c(5)]
names(adultsample)[-c(1,6,8,9)]
names(synadult\$syn)
##------ native.country moved to end as many levels that slow down synthesis-----
synadult <- syn(adultsample, visit.sequence = c(1:12,14,13) ,seed = 7868)
##------ get basic disclosure and make Figure 2----------------------------------
dd1 <- multi.disclosure(synadult,adultsample,
   keys = c("age" , "occupation","race", "sex" ))
pdf("vignettes/fig2dis.pdf",height=300, width=500)
print(dd1)
dev.off()
##------ exclude some target levels and missing values for targets---------------
dd2 <- multi.disclosure(synadult,adultsample,
    print.flag = FALSE, not.targetslev = c(rep("",5),"0","0","","United-States",""),
    keys = c("age" , "occupation","race", "sex" ))
##------ exclude some target levels and missing values for targets---------------
dd3 <- multi.disclosure(synadult,adultsample, print.flag = FALSE,
   not.targetslev = c(rep("",5),"0","0","","United-States",""),
    keys = c("age" , "occupation","race", "sex" ))
    dd1 <- multi.disclosure(synadult,adultsample, keys = c("age" , "occupation","race", "sex" ))
##-------restrict to denominators of 1 -----------------------------------------
dd4 <- multi.disclosure(synadult,adultsample, denom_lim =1,
    not.targetslev = c(rep("",5),"0","0","","United-States",""),
    usekeysNA = FALSE, usetargetsNA = FALSE,
    exclude_ov_denom_lim = TRUE,
    keys = c("age" , "occupation","race", "sex" ))
##----------just denominators not others---------------------------------------
dd5 <- multi.disclosure(synadult,adultsample, denom_lim =1,
    exclude_ov_denom_lim = TRUE,
    keys = c("age" , "occupation","race", "sex" ))
##-----------make table 1------------------------------------------------------
library(stringr)
cols12 <- dd1\$attrib.table[,1:2]
  dimnames(cols12)[[1]] <- word(dimnames(cols12)[[1]],2,2)
cols34 <- dd2\$attrib.table[,1:2]
  dimnames(cols34)[[1]] <- word(dimnames(cols34)[[1]],2,2)
cols56 <- dd3\$attrib.table[,1:2]
  dimnames(cols56)[[1]] <- word(dimnames(cols56)[[1]],2,2)
cols78 <- dd4\$attrib.table[,1:2]
  dimnames(cols78)[[1]] <- word(dimnames(cols78)[[1]],2,2)
cols910 <- dd5\$attrib.table[,1:2]
  dimnames(cols910)[[1]] <- word(dimnames(cols910)[[1]],2,2)
table.adult <- data.frame(cols12[order(dimnames(cols12)[[1]]),],
    cols34[order(dimnames(cols34)[[1]]),],
    cols56[order(dimnames(cols56)[[1]]),],
    cols78[order(dimnames(cols78)[[1]]),],
    cols910[order(dimnames(cols910)[[1]]),])
table.adult
##--------------------- check 2way for one target------------------------------
dis1 <- disclosure(synadult,adultsample, target = "workclass",
    keys = c("age" , "occupation","race", "sex" ))
print(dis1, to.print = c("attrib","check\_2way"))

### v4

@@
 Abstract
-This paper explains how synthpop package111see www.synthpop.org.uk  [^2] The new version 1.9-0 is currently available on Github at https://github.com/gillian-raab/synthpop has been extended to include functions to calculate identity and attribute disclosure risk measures for the original records.
-using information obtained from the corresponding synthetic data.
+This paper explains how synthpop package 111see www.synthpop.org.uk  [^2] The new version 1.9-3 is currently available on Github at https://github.com/gillian-raab/synthpop has been extended to include functions to calculate measures of identity and attribute disclosure risk for synthetic data that measure risks for the records used to create the synthetic data.
+This paper is also appears as a vignette as pat of the versions of the package from 9.3 onwards.
 The basic function disclosure calculates identity disclosure for a set of quasi-identifiers (keys) and attribute disclosure for one variable specified as a target from the same set of keys.
-The second function multi.disclosure is a wrapper for the first and it presents summary results for a set of targets.
+The second function multi.disclosure is a wrapper for the first and presents summary results for a set of targets.
 This short paper explains the measures of disclosure risk and documents how they are calculated.
-We recommend two measures: $RepU$ (replicated uniques) for identity disclosure and $DiSCO$ (disclosive in synthetic and correct in the original) for attribute disclosure.
-Both are expressed a % of the original records and both can be compared to similar measures calculated from the original data.
+We recommend two measures: $RepU$ (replicated uniques) for identity disclosure and $DiSCO$ (Disclosive in Synthetic Correct Original) for attribute disclosure.
+Both are expressed a % of the original records and each can be compared to similar measures calculated from the original data.
 Experience with using the functions on real data found that some apparent disclosures could be identified as coming from relationships in the data that would be expected to be known to anyone familiar with its features.
@@
 1 Introduction
-In his recent review of thirty years of synthetic data Reiter [17] comments:
+In his recent review of synthetic data (SD) methodology Reiter [18] comments:
 ”While there is need to examine disclosure risks in synthetic data, there is no standard for doing so, especially in fully synthetic data. Instead, disclosure checks tend to be ad hoc”
-This is in contrast to the variety of measures of utility available for synthetic data; see [16] for a list of these.
-While utility measures must be chosen that are relevant to the intended use of the synthetic data, disclosure measures must focus on the possible harm to the privacy of an individual or other unit whose data contributed to the creation of the synthetic data.
-The evaluation of the disclosure risk from synthetic data must relate to the context of its release (see [5] for a discussion of this).
+This is in contrast to the variety of measures of utility available for SD; see [15] for a list of these.
+While utility measures must be chosen that are relevant to the intended use of the SD, disclosure measures must focus on the possible harm to the privacy of an individual or other unit whose data contributed to the creation of the SD.
+The evaluation of the disclosure risk from SD must relate to the context of its release (see [5] for a discussion of this).
 We cannot expect a fixed rule, for example that a criterion for release requires a value of some disclosure measure beyond a threshold.
-Instead, we expect those releasing data to use the disclosure measures to evaluate potential harm to the data subjects or to the data custodians from information in the released data.
-We hope that these new functions will allow the disclosure risk of synthetic data sets to be explored and, where necessary, reduced.
-We expect that disclosure risks from synthetic data to be lower than those from the original data.
+Instead, we expect those releasing data to use the disclosure measures to evaluate potential harm to the data subjects, or to the data custodians, from information in the released data.
+We hope that these new functions will allow the disclosure risk of SD sets to be explored and, where necessary, reduced.
+We expect that disclosure risks from SD to be lower than those from the original data.
 But in some cases, e.g.
 a large data set with only a few categorical variables, the disclosure risk from the original may already be low.
-To evaluate the disclosure risk of a synthesis method, the risks for synthetic data must be compared with equivalent risks for the original.
+To evaluate the protection from disclosure afforded by a synthesis method, the risks for SD must be compared with equivalent risks for the GT.
 We consider two types of disclosure risk:
@@
 This is in contrast to incomplete or partial synthesis where only some variables are replaced.
-data because there is no one-to-one correspondence between records in the original and synthetic data sets.
+data because there is no one-to-one correspondence between records in the original and SD.
 But it may still be of interest since it is an important factor for attribute disclosure.
   • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the original data.
-The measures we describe here are appropriate for fully synthetic data where all items in all records are replaced by synthetic values.
-Different measures have been developed for partially synthetic data [3, 18]
-The disclosure risk posed by synthetic data can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques.
-These methods can be used to reduce disclosure risk by modifying the original data before synthesis, or the synthetic data before its is released.
+The measures we describe here are appropriate for fully SD where all items in all records are replaced by synthetic values.
+Different measures have been developed for partially SD [3, 19].
+Note that our disclosure methods treat numeric variables, by default, as if they were categories unless ngroups_keys and/or ngroups_target(s) are set.
+In our first example the income variable has been grouped into 20 categories, but the other numerical variable (depress) with only 21 categories has been left as it is.
+The number of groups formed may differ from the parameter setting e.g.
+if there are fewer than numgroups distinct values.
+[^4] The code uses grouping options from functions in the classInt package..
+The disclosure risk posed by SD can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques.
+These methods can be used to reduce disclosure risk by modifying the original data before synthesis, or the SD before its is released.
 Some such methods are already available in the synthpop package.
 These include categorising, top/bottom coding and smoothing for continuous variables, and the merging of small categories for factors.
-The removal of replicated uniques is another option available.
-There has been many recent proposals for making synthetic data sets Differential Private (DP) (selection of refs).
+The removal of replicated uniques is another option available in synthpop.
+There have been many recent proposals for making synthetic data sets comply with Differential Privacy (DP) [4].
 DP is a very strong privacy guarantee that protects against an intruder with arbitrary external information about the subjects in the data, except for the one whose privacy is being protected.
-This is an unrealistic assumption and DP synthetic data has been shown to have unacceptably low utility in many cases [1, 20].
-We will not discuss these methods here, but note that we could use the metrics proposed here to evaluate disclosure risks for DP synthetic data.
+This is an unrealistic assumption and DP SD has been shown to have unacceptably low utility in many cases [1, 21].
+We will not discuss these methods here, but note that we could use the metrics proposed here to evaluate disclosure risks for DP SD.
 2 A simple example
 Here we illustrate the basic use of multi.disclosure.
-If the parameter targets is not specified, all the variables in the synthetic data that are not part of keys are used as targets.
+If the parameter targets is not specified, all the variables in the SD that are not part of keys are used as targets.
 The identity disclosure measures are UiO for original and repU for synthetic, and for attribute disclosure Dorig for original and DiSCO for synthetic.
 These and other measures will be explained in Sections 3.2 and 3.3.
-First, a subset of 9 variables are selected fron the SD2011 data (a survey on quality of life in Poland) that is available as part of the synthpop package.
+First, a subset of 9 variables are selected from the SD2011 data (a survey on quality of life in Poland) that is available as part of the synthpop package.
 A single synthetic data set is created by the default method in synthpop: cart for each conditional distribution.
-The synthetic data object s1444an object of class synds has a component syn that is a single synthetic data set.
+Note that the variable income has values -8 that indicate not applicable, and the synthesis allows for this.
+The synthetic data object s1555an object of class synds has a component syn that is a single synthetic data set.
 The disclosure functions can also be used with synthetic data created by other methods either as single synthetic data sets or lists of repeated syntheses from the same original.
 Here we select 4 keys that represent items that might be known for members of this sample, or of the Polish population in 2011.
-By default,the 5 other variables become the targets: income (rounded to give 406 distinct values) marital (marital status), workab (intention to work abroad), ls and depress (scores for life satisfaction and depression with 7 and 21 distinct values).
-We first use multi.disclosure to create a multi.disclosure object for these keys, and we print out its identity component.
+By default,the 5 other variables become the targets: depress income ls marital (marital status), workab (intention to work abroad).
+The second target, income, is grouped with a target of 20 categories [^6] The algorithm attempts to create this number of groups with approximately equal sizes.
+Skew data often result in fewer groups.
+In this case 18 groups are formed from the numeric values, plus two groups from the -8 code and the missing values, giving a total of 20 We first use multi.disclosure to create a multi.disclosure object for these keys, and we print out its identity component.
+$\begin{Schunk}$
+$\begin{Sinput}$
 R> library("synthpop")
@@
 +    "income","ls","marital" , "workab")]
-R> s1 <- syn(ods, seed = 8564, print.flag = FALSE)
+R> s1 <- syn(ods, seed = 8564, print.flag = FALSE, cont.na = list(income = -8))
 R> t1 <- multi.disclosure(s1, ods, print.flag = FALSE, plot = FALSE,
-+     keys = c("sex", "age", "region", "placesize"))
++  keys = c("sex", "age", "region", "placesize"),   ngroups_targets = c(0,20,0,0,0))
 R> print(t1,to.print = "ident")
+[…]
+$\begin{Soutput}$
 Disclosure risk for 5000 records in the original data
@@
 For synthetic ( repU ) 14.86 %.
+[…]
+[…]
 The measure $UiO$ (Unique in Original) shows that 48% of the original records would have unique combinations of these 4 keys.
@@
 Details of how identity and disclosure measures are calculated can be found in Section 3.
+The results in the Figure 1 are in descending order by the attribute disclosure risk in the SD.
 Figure 1: Plot from multi.disclosure with plot = TRUE, the default value.
-The results in the plot are in descending order by the disclosure in the synthetic data.
 The measure $Dorig$ tells us that these 4 keys would identify a unique value of each of these targets for over 50% of the original records.
-$DiSCO$, the proportion of the original records that are disclosive in the original and also in the synthetic data with a correct attribution to the target give lower values.
+We get lower values for $DiSCO$, the proportion of the original records that are disclosive in the original and also in the synthetic data with a correct attribution to the target.
 The first two variables shown in Figure 1 have additional labels that flag possible contributions to disclosure from knowledge of 1-way or 2-way relationships in the original data.
 This is detailed in Section 4.
-Previous work on attribute disclosure [4, 22] has used the Correct Attribution Probability ($DCAP$) as a disclosure measure for synthetic data.
-This is calculated as the percentage of records correctly predicted from the synethetic data by a set of keys.
-However it has been suggested [7, 2], that it may be as much a measure of utility as disclosure risk and modifications have been suggested [21, 10] one of which ($TCAP$) is close to our proposed measure, $DiSCO$.
-For completeness the measures $DCAP$ and $TCAP$ are calculated by disclosure; see Appendix 2.
 3 Scenario and definitions
@@
 ”Fake Data” are thought just to be a cover up.
+It may also be a reqasonable measure to use without this scenario, since it compares the disclosiveness of the SD to that of the GT.
 Disclosure measures from the synthetic data are each compared to similar measures for someone with access to the original data.
@@
 First proposed in 1998 [20] it is discussed fully in [5].
-A table is k-anonymous if a set of keys identifies at most $k-1$ individuals.
+A table is k-anonymous if a set of keys identifies at most $k$ individuals.
 Thus 2-anonymous data will never identify just one individual.
 Based on this idea, the percentage of records for which the keys identify just one individual give identity disclosure measures.
-A table of $q$ values is produced from the synthetic and the original data.
-$UiO$ and $UiS$ are the percentages of records with keys where the table value is 1.
+Tables of $q$ values are produced from the synthetic and the original data.
+$UiO$ and $UiS$ are the percentages of records with keys where the table count is 1.
 An intruder checking out a record for their known set of keys will look for it in the synthetic data.
-Some records will not be in the synthetic data and $UiOiS$ (Unique in Original in Synthetic) gives the % that would be found.
-These records are then checked for uniqueness in the synethetic data giving, $repU$ is the percentage of unique original records that are also unique in the synthetic data.
-The percentage $repU$ has been used as a disclosure measure to evaluate synthetic data by [9] and by [15]  [^7] Jackson et al.
+Some records will not be in the SD and $UiOiS$ (Unique in Original in Synthetic) gives the % that would be found.
+These records are then checked for uniqueness in the synthetic data giving, $repU$ is the percentage of unique original records that are also unique in the SD.
+The percentage $repU$ has been used as a disclosure measure to evaluate SD by [9] and by [17]  [^9] Jackson et al.
 in [9] argue that the denominator for $repU$ should be $N_{s}$ rather than $N_{d}$.
 This is inappropriate because our scenario is to consider the risk to the original data..
-Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the synthetic data.
+Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the SD.
 Nowok et al.
 [14] have evaluated this and give an example where this process has very little effect on utility.
-The function replicated.uniques [^8] For example by replicated.uniques (s2, ods, keys = c("sex","region","age","placesize") also calculates $repU$ using a different method from the one described here.
-One of the outputs of the functions disclosure and multi.disclosure is ident, a table of identity disclosure measures as illustrated by this example, using the keys from the example in Section 2 but now calculated for a synthetic object with 5 data sets and using the one target depress.
+The function replicated.uniques [^10] For example by replicated.uniques(s2, ods, keys = c("sex","region","age","placesize")) also calculates $repU$ using a different method from the one described here.
+One of the outputs of the functions disclosure and multi.disclosure is ident, a table of identity disclosure measures as illustrated by this example, using the keys from the example in Section 2 but now calculated for a synthetic object with 5 data sets, using the one target depress.
 We first create t5 an object of class disclosure and then print out the identity and attribute disclosure measures for each synthetic data set.
 The table for identity disclosure has $UiO$ and $repU$ as its first and last column.
-The 2nd and 3rd columns are $UiS$ calculated from the synthetic data in the same manner as $UiO$ from the original and $UiOiS$ the % of $UiO$ with $q$ that are in the synthetic data but not necessarily unique.
+The 2nd and 3rd columns are $UiS$ calculated from the synthetic data in the same manner as $UiO$ from the original and $UiOiS$ the % of $UiO$ with $q$ that are in the SD but not necessarily unique.
 Each are steps towards calculating $repU$.
+$\begin{Schunk}$
+$\begin{Sinput}$
 R> s5 <- syn(ods, seed = 8564, m = 5, print.flag = FALSE)
@@
 R> print(t5, to.print = c("ident"))
+[…]
+$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
@@
 5 48.38 36.14 22.00 14.62
+[…]
+[…]
 3.3 Attribute disclosure measures
-We approach disclosure risk from the point of view of an intruder with access to the synthetic data and to certain attributes (quasi-identifiers) known for one or more individuals in the original data.
-The quasi-identifiers for the known record(s) can be combined to create a composite variable $q$ that can be created for all records in the original and synthetic data.
-Modelling what an intruder might do we calculate the following measures, each of whuch is a proportion of the original records:
-  • Look up $q$ for each original record in the synthetic data. The proportion found becomes $iS$ (in Synthetic).
+The attribute disclosure measure for the SD that is displayed in Figure 1 is our preferred measure, $DiSCO$ [16].
+In section 3.3.1 we explain it via the example above.
+Other related disclosure measures, $DCAP$ and $TCAP$, have been proposed [6, 23, 22, 11, 10].
+The disclosure function also computes these measures.
+Section 3.3.2 describes them and discusses circumstances where they differ from $DiSCO$.
+The reason for our preference for $DiSCO$ compared to other proposed methods is that it is a mesure computed for all the records in the original data.
+The other measures thus do not include aspects of disclosure, for example that some records in the original not appearing in the SD, that are valid contributions to the disclosure reduction effected by the synthesis.
+The measure we describe as $DCAP$ is readily adapted to apply to all original records by replacing $N_{s}$ in its denominator with $N_{d}$.
+When the synthesis creates SD of the same size as the original $DCAP$ and $Dorig$ are identical.
+But, as we show in Figure 2, $DCAPorig$ assesses disclosure appropriately when synthetic data with fewer records than the original are created.
+This measure is added to the component attribattrib of the disclosure object created by the disclosure function.
+The two measures $DiSCO$ and $DCAPorig$ can be considered as the extremities of a range of a more general disclosure measure when a record is considered disclosive if its probability of correct attribution reaches a threshold tautau.
+DiSCODiSCO corresponds to a tautau of 1.0 while $DCAPorig$ corresponds to a tautau of zero.
+This more general approach may be included in future releases of synthpopsynthpop.
+Appendix 1 gives formal definitions of all disclosure measures discussed here.
+3.3.1 Our preferred measure: Disclosive in Synthetic Correct in Original ($DiSCO$)
+We approach disclosure risk from the point of view of an intruder with access to the SD and to certain attributes (quasi-identifiers) known for one or more individuals in the original data.
+The quasi-identifiers for the known record(s) can be combined to create a composite variable $q$ that can be created for all records in the original and SD.
+Modelling what an intruder might do we calculate the following measures, each of which is a proportion of the original records:
+  • Look up $q$ for each original record in the SD. The proportion found becomes $iS$ (in Synthetic).
   • Check if all records with the same $q$ have the same level of the target $t$. The proportion passing this further test becomes $DiS$ (Disclosive in Synthetic).
-  • Then check if these apparent disclosures corresponds to the value of $t$ in the original data.
-The proportion of original records for which this is true becomes $DiSCO$ Disclosive in Synthetic Correct in Original.
+  • Then check if these apparent disclosures correspond to the value of $t$ in the original data.
+The proportion of original records for which this is true becomes $DiSCO$ Disclosive in Synthetic Correct Original.
 Note that records contributing to $DiSCO$ may not be disclosive in the original data, as this information would not be available to the intruder.
 A further measure $DiSDiO$ (Disclosive in Synthetic Disclosive in Original) restricts the score to those disclosive in the original.
-These measures are defined formally in Appendix 1.
+$\begin{Schunk}$
+$\begin{Sinput}$
 R> print(t5, to.print = c("attrib"))
+[…]
+$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
 Attribute disclosure measures for depress from keys: sex age region placesize
-  Dorig  Dsyn    iS   DiS DiSCO DiSDiO max_denom mean_denom
-1  53.3 46.26 64.90 34.18  9.54   6.14         3       1.16
-2  53.3 44.80 64.00 32.50 10.26   6.78         4       1.19
-3  53.3 44.60 64.02 32.14  9.10   5.92         4       1.19
-4  53.3 45.80 63.88 33.38  9.20   5.52         4       1.21
-5  53.3 44.52 63.44 31.46  9.34   5.80         4       1.23
+  Dorig  Dsyn    iS   DiS DiSCO DCAPorig DiSDiO max_denom mean_denom
+1  53.3 46.26 64.90 34.18  9.54    16.39   6.14         3       1.16
+2  53.3 44.80 64.00 32.50 10.26    17.45   6.78         4       1.19
+3  53.3 44.60 64.02 32.14  9.10    16.20   5.92         4       1.19
+4  53.3 45.80 63.88 33.38  9.20    15.92   5.52         4       1.21
+5  53.3 44.52 63.44 31.46  9.34    16.17   5.80         4       1.23
+[…]
+[…]
 Here we print the table of attribute disclosure measures for the example in the previous section.
 As we move from $iS$ to $DiS$ and to $DiSCO$ we can see how the %disclosive is affected by different conditions.
-Here lack of $q$ levels in the synthetic data retains just 64% of records.
-The requirement for a record to be disclosive in the synthetic data reduces this to 34% and again to around 9% for those with a corect attribution.
+Here lack of $q$ levels in the SD retains just 64% of records.
+The requirement for a record to be disclosive in the SD reduces this to 34% and again to around 9% for those with a correct attribution.
 This reduces it to around 6% by restricting to records disclosive in the original.
-The $Dorig$ and $DiSCO$ measures are not restricted to disclosures that are identified from unique records for $q$ in either the original or the synthetic data.
-The number of records contributing to each disclosive $qt$ cell in the original table is the denominator that applies to that record.
-The columns max denom and mean denom for the records disclosive in the synthetic data contributing to the $DiSCO$ refer to the denominators ($d_{.q}$ used to calculate $pd_{tq}$ from the original data).
-We can see from the mean that here the majority of disclosive records had unique key combinations in the original data, and the maxima was 3 for the first synthesis and 4 for the others.
+The $Dorig$ and $DiSCO$ measures are not restricted to disclosures that are identified from unique records for $q$ in either the original or the SD.
+The number of records contributing to each disclosive $qt$ cell in the synthetic table is the denominator that applies to that record.
+The columns max denom and mean denom refer to the denominators for the records disclosive in the original data that contribute to $DiSCO$.
+We can see from the mean that here the majority of disclosive records had unique key combinations in the SD, and the maxima was 3 for the first synthesis and 4 for the others.
 Large denominators can be an indication that some of the disclosure may be coming from strong relationships between variables in the data that might even be expected a-priori.
@@
 The example in Section 2 is here run restricted to denominators of 1.
+$\begin{Schunk}$
+$\begin{Sinput}$
 R> multi.disclosure(s1, ods, print.flag = FALSE, plot = FALSE,
@@
 +    denom_lim =1, exclude_ov_denom_lim = TRUE)
+[…]
+$\begin{Soutput}$
 Disclosure risk for 5000 records in the original data
@@
 Variables Ordered by synthetic disclosure measure
-                     attrib.orig attrib.syn check1 Npairs check2
-1 income                   48.38       4.02             0
+                     attrib.orig attrib.syn                   check1 Npairs
+1 income                   48.38       3.36                               0
 2 depress                  48.38       7.02                               0
-3 ls                       48.38       9.74             0
-4 marital                  48.38      15.72             0
-5 workab 1way checks       48.38      20.62     NO      0
+3 ls                       48.38       9.84                               0
+4 marital                  48.38      15.40                               0
+5 workab 1way checks       48.38      20.48 Check  workab  level  NO      0
+                     check2
+1 income
+2 depress
+3 ls
+4 marital
+5 workab 1way checks
+[…]
+[…]
 The $DiSCO$ values have decreased, as expected, but $UiO$ and $repU$ have increased and also that $Dorig$ is now the same for all targets and equal to the $UiO$ value before the denominator exclusion.
@@
 This and other methods of excluding certain apparent disclosures are discussed in the next section.
+3.3.2 Other attribute disclosure measures
+Previous work on attribute disclosure [6, 23] has used the Correct Attribution Probability ($DCAP$) as a disclosure measure for synthetic data.
+This is calculated as the average percentage of records correctly predicted from the synthetic data by a set of keys.
+Two different versions are described to deal with the case when some key combinations in the SD do not occur in the original, so that the attribution probability is undefined.
+In one case, that we will designate as $DCAP$, the undefined attribution probabilities are set to zero and the denominator is the total number of synthetic records.
+The second case $DCAP_{alt}$ [^11] This approach is sometimes referred to as $TCAP_{undefined}$ [23].
+uses the number of synthetic records present in the synthetic data as the denominator, so that $DCAP_{alt}\geq DCAP$.
+The related quantity $TCAP$ restricts the probabilities to those over a certain limit, and most commonly this limit is $1.0$ making the numerator of the $TCAP$ measures identical to that for $DiSCO$.
+One possibility for a denominator would be to use just the number of synthetic records with key combinations in the original, as for $DCAP_{alt}$ [^12] This measure was what was reported as $TCAP$ in Versions of $Synthpop$ before Version 1.9.3.
+But evaluations using $TCAP$ [11, 10] use an even smaller denominator; the number of synthetic records with keys present in the original and where the target probaility is 1.0.
+This can be described as the proportion correctly predicted for all apparently disclosive records; we refer to this measure as $TCAP$ and note that its smaller denominator implies that $TCAP\geq TCAP_{alt}\geq DiSCO$ when the synthetic data set(s) are the same size as the original.
+All of these measures are shown in Table 1 for 3 of the targets in the synthesis of the 5,000 records from the example above.
+We can see that the inequalities above all hold.
+For this case the ordering of the disclosive risk by target is consistent for all measures.
+For the three measures that share the same numerator ($DiSCO$, $TCAP_{alt}$ and $DCAP$) we see that the reduced denominators have a large effect in this example, so that $TCAP$ is around three times $DiSCO$.
+DiSCO DCAP DCAPalt TCAPalt TCAP
+ls 12.63 21.97 29.57 17.00 36.71
+depress 9.21 16.23 21.85 12.40 28.57
+income 6.16 11.34 15.27 8.29 19.96
+Table 1: Different disclosure measures (%) for 3 targets from the example above when number of synthetic records equals the number of original records (5000).
+Another way thatt he disclosure risk to the original records can be reduced is to release synthetic data sets that are smaller than the original.
+This is the approach that has been used when UK Census data has been released [2] or for low-fidelity Census data from Scotland [^13] https://www.researchdata.scot/news-and-insights/synthetic-census-data-now-available-for-research/ accessed 7/5/25.
+Figure 2 illustrates how the different disclosure measures assess the risks in this example for sizes of SD from 5000, as the original, to only 250 records.
+Of the measures in Table 1, only $DiSCO$ and $DCAPorig$ adjust the attribute disclosure risk for the reduced number of records in the SD.
+The $TCAP$ measure is almost independent of the size of the SD.
+The other measures all give misleadingly high disclosure risks for small samples.
+Figure 2: Plot of disclosure measures for synthetic data of size ($k$) from 250 to 5000.
+Figure 2 displays
 4 Identifying disclosure from 1-way and 2-way relationships
@@
 The first is a check for a target where a high proportion of records have one level of the target.
-The second is when there is a strong relationship between the target and one of the keys, so that one $tq$ pair accounts for many of the disclosive records.
-These two aspects are flagged by the values of check_1way and check_2way that are returned as part of objects of Returning to the example in 2 we now use the function disclosure to get details for the two targets that were flagged as requiring checking; see Figure 1.
+The second is when there is a strong relationship between a target and one of the keys, so that one $tq$ pair accounts for many of the disclosive records.
+These two aspects are flagged by the values of check_1way and check_2way that are returned as part of objects of Returning to the example in Section 2, we now use the function disclosure to get details for the two targets that were flagged as requiring checking; see Figure 1.
+$\begin{Schunk}$
+$\begin{Sinput}$
 R> d1_workab <- disclosure(s1, ods, print.flag = FALSE, target = "workab",
 +     keys = c("sex", "age", "region", "placesize"),plot = FALSE)
-R> print(d1_workab, to.print = c("check_1way","check_2way"))
+R> print(d1_workab, to.print = c("check_1way"))
+[…]
+$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
 Details of target level contributing disproportionately to disclosure
-  check Level totalDisclosive PctDisAll totalLevel pctdisLevel
-1     1    NO            2661     53.22       2529       95.04
-Details of target-key combinations contributing disproportionately to disclosure
-This is a list with one for each of syntheses
-[[1]]
-  total pctall max_denom    key        target_key pct_2way denom_2way
-1     6    0.2         6    age           NO | 55     97.1        105
-2     6    0.2         6 region NO | Podkarpackie     95.2        313
+   Level  All PctLevelAll totalDisclosive nLevelDis PctLevelDis
+NO    NO 5000       88.64            2605      2482       95.28
+Details of target-key pairs contributing disproportionately to disclosure of workab
+26 pairs need checks
+Please examine component \$check_2way of the disclosure object and look at original data.
+Consider excluding these key-target pairs with some the following parameters to disclosure:
+exclude_ov_denom_lim = TRUE or defining key-target combinations from exclude.targetlevs,
+exclude.keys and exclude.keylevs
+[…]
+[…]
 We can see that it was the category ”NO” of workab that contributed most to the disclosure risk; most survey respondents (89%) had never worked abroad.
-This level of the target accounted for 93% of apparent disclosures for workab.
+This level of the target accounted for 95% of apparent disclosures for workab.
 To predict this level for all of a group with the same $q$ could hardly be considered disclosive.
-This would be true if the intruder had access to the marginal distribution of workab, and even if they did not some knowledge of the respondents’ background might suffice.
+This would be true if the intruder had access to the marginal distribution of workab, and even if they did not, some knowledge of the respondents’ background might suffice.
 The $tq$ pairs identified for workab both included the ”NO” level of the target.
-[…]
+$\begin{Schunk}$
+$\begin{Sinput}$
 R> d1_marital <- disclosure(s1, ods, print.flag = FALSE, target = "marital",
 +     keys = c("sex", "age", "region", "placesize"),plot = FALSE)
-R> print(d1_marital, to.print = c("check_1way","check_2way"))
+R> print(d1_marital, to.print = c("check_2way"))
+[…]
+$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
-No 1 way checks for your target are flagged at your settings for thresh_1way
 Details of target-key combinations contributing disproportionately to disclosure
-This is a list with one for each of syntheses
-[[1]]
-  total pctall max_denom key   target_key pct_2way denom_2way
-1     7    0.4         7 age MARRIED | 40     82.3         79
-2     6    0.3         6 age  SINGLE | 20    100.0         79
+This is a list with one for each of 4 6 syntheses
+  target_key_levs npairs key key_target_total key_total PctTargetKeyLevel
+7       SINGLE|19     10 age               91        92             98.91
+4      MARRIED|41      5 age               59        73             80.82
+6       SINGLE|18      5 age               91        92             98.91
+8       SINGLE|22      5 age               85        91             93.41
+[…]
+[…]
 For the target marital two $tq$ pairs are identified as contributing large denominators.
-Those aged 20 are all SINGLE, and those aged 40 are mainly MARRIED.
+Those aged 19 are almost all SINGLE, and those aged 40 are mainly MARRIED.
 Again these would not be considered disclosive for an intruder with some knowledge of the respondents.
@@
 For one-way disclosure the parameter thresh_1way has the default value of $c(50,90)$, meaning that there must be at least 50 disclosive records for one level of a target and that $q$ values including this target must account for over 90% of all disclosive records.
-For two way relationships the thresh_2way has default value $c(5,80)$ and the algorithm first identifies all disclosive $tq$ combinations with denominators over 5.
+For two way relationships the thresh_2way has default value $c(5,80)$ and the algorithm first identifies all disclosive $tq$ combinations with denominators over 4.
 It then identifies the level of the key in $q$ with that best predicts this level of $t$ and checks if over 80% of the disclosive records with this key would have the correct prediction.
 5 Excluding records
-Having identified records where the disclosure is not really telling us anything new, one option is to exclude some of these original records explicitly from the measures.
+Having identified records where the apparent disclosure is something that would generally be known, one option is to exclude these key-target combinations explicitly from the measures.
 The following parameters for multi.disclosure and disclosure can be used for this.
@@
   • usekeysNA:This is set to TRUE by default so that missing values are included in all tables. It can be set to FALSE for some or all keys to exclude NAs.
-  • usetargetsNA)(multi.disclosure) or usetargetNA (disclosure) ): Similar to the above for target(s) - note can be a vector for multi.disclosure.
-  • exclude.keys, exclude.keylevs and exclude.targetlevFor disclosure only: Three vectors of length the number of key-target pairs to be excluded from all tables.
+  • usetargetsNA (multi.disclosure) or usetargetNA (disclosure): Similar to the above for target(s) - note can be a vector for multi.disclosure.
+  • exclude.keys, exclude.keylevs and exclude.targetlev: Three vectors of length the number of key-target pairs to be excluded from all tables.
+[^14] For multi.disclosure these parameters are supplied as lists with elements for each target in targets
 Figure 3: Plot from multi.disclosure for variables flagged by checks.
 To illustrate exclusions we will use the Adult data set from the UCI machine learning repository [13] with almost 50 thousand records from the US Census income study, available as part of the arules package for R [8].
-This is one of the data sets used in [6] to evaluate privacy risks for synthetic data.
-Figure 2 is the output of multi.disclosure from the keys age, occupation, race and sex for the other 10 variables in the data.
-The code used to calculate the results in this section is in Appendix 3.
+This is one of the data sets used in [7] to evaluate privacy risks for SD.
+Note that the two variables, capital gain and texttcapital gain, originally numeric data, have been truncated so that they have only 123 and 99 distinct values repsectively.
+They have not been grouped further in the results presented here.
+Figure 3 is the output of multi.disclosure from the keys age, occupation, race and sex for the other 10 variables that are not in the keys.
 The disclosiveness of the original data is relatievly low, compared to the previous example, because of the large sample size and the absence of any geographic identifiers.
 Three of the 10 variables capital.gain, capital.loss and native.country are flagged to check one-way relationships.
-The check1 column of the attribute table tells us that the levels contributing to disclosre are zeros for the first two and United-States for native.country.
+The check1 column of the attribute table tells us that the levels contributing to disclosure are zeros for the first two and United-States for native.country.
 These levels make up 95%, 92% and 89% of all records.
@@
 The summary function multi.disclosure does not allow target and key specific pairs to be excluded.
-To investigate this it is necessary to examine the output of disclosure999see code in Appendix 3 for how to do this..
+To investigate this it is necessary to examine the output of disclosure151515see code in Appendix 3 for how to do this..
 This showed that the largest contribution to check_2way for workclass was due to this being missing when occupation was missing, although other relationships between these two variables also contributed.
@@
 Table 2: Disclosure results from Adult data with different exclusions.
-Table 1 gives the results of excluding different entries in the tables of $q$ and $t$ from the attrinute disclosure measures.
-Excluding the levels of the target flagged by textttcheck_1way reduces the $DiSCO$ to almost zero for these 3 variables.
+Table 2 gives the results of excluding different entries in the tables of $q$ and $t$ from the attribute disclosure measures.
+Excluding the levels of the target flagged by check_1way reduces the $DiSCO$ to almost zero for these 3 variables.
 Adding exclusion of missing values reduced the disclosure for workclass and for some other variables a little.
@@
 DP claims to protect data from an intruder with arbitrary knowledge of the data, except of the one record that has the greatest influence on the likelihood of the results.
-In contrast our metrics require the user to specify the details of what they would expect an intruder to know about the data.
-Most importantly, the user must specify keys that identify variables in the data that would expect to be known about individuals.
+In contrast our metrics require the user to specify keys that identify variables in the data that would expect to be known about individuals, as well as to specify the details of what they would expect an intruder to know about the data.
 Also, the routines flag cases where part of the disclosure measures come from one- or two-way relationships so that disclosure would be expected even in the absence of data.
 We have evaluated the routines on a few data sets and set levels of the thresholds for this at what seem to be reasonable levels, but more experience with other data sets would be valuable.
-We hope that these tools will be helpful to data holders who need to make decisions about the risks of releasing synthetic data to the public, or to a restricted audience.
+We hope that these tools will be helpful to data holders who need to make decisions about the risks of releasing SD to the public, or to a restricted audience.
 They should also enable the disclosure risks of different synthesis methods to be evaluated.
 7 Acknowledgement
-Research Data Scotland (https://www.researchdata.scot/ funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package101010 from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control.
+Research Data Scotland (https://www.researchdata.scot/) funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package161616from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control.
 We also thank the Scottish Centre for Administrative Data Research for continue to support the development of the synthpop package since its creation was supported by the ESRC funded SYLLS poject in 2012-14.
-We would also like to thank Timon Huijser for noting some mistakes in an earlier version of this paper that are now corrected
+We would also like to thank Timon Huijser, for pointing out errros in a previous version of this vignette.
 Appendix 1: Notation and formal definitions
 Before defining the measures of identity and disclosure risk we need to introduce the notation that will be used to calculate them.
-The first step is to create the quasi-identifiers from the keys for the original and synthetic data.
+The first step is to create the quasi-identifiers from the keys for the original and SD.
 For the keys used in the example given in Section 2 the quasi-identifier that we will designate as $q$ for the first record in the original data is:
 "FEMALE | 57 | Lubuskie | URBAN 100,000-200,000"
-and that for the first record in the synthetic data:
+and that for the first record in the SD:
 "FEMALE | 39 | Zachodnio-pomorskie | URBAN 100,000-200,000".
-In order to calculate identity disclosure measures, we need to compare the tables of $q$ from the original and synthetic data.
-For attribute disclosure measures we need to cross-tabulate $q$ with each target variable $t$ and compare findings from the synthetic data with what would have been found from the original data.
-In general, the levels of $q$ and sometimes $t$ in the original and synthetic data will not be the same.
+In order to calculate identity disclosure measures, we need to compare the tables of $q$ from the original and SD.
+For attribute disclosure measures we need to cross-tabulate $q$ with each target variable $t$ and compare findings from the SD with what would have been found from the original data.
+In general, the levels of $q$ and sometimes $t$ in the original and SD will not be the same.
 Before creating any tables, we need to define sets of $q$ and $t$ values that give the union of both sets of levels and align the tables so that their indices correspond.
-For the original data $d_{.q}$ is the count of records with the keys corresponding to the levels of $q$ and $q_{tq}$ the count of records with this $q$ and level $t=1,...T$ of the target.
+For the original data $d_{.q}$ is the count of records with the keys corresponding to the levels of $q$ and $q_{tq}$ is the count of records with this $q$ and level t=1,Tt=1,...T of the target.
 The equivalent counts from the synthesised data are designated by $s_{.q}$ and $s_{tq}$.
 When a member of $q$ is in the original data but not in the synthetic, $s_{.q}$ and $s_{tq}$ are all zero.
-Similarly when a member of $q$ is in the synthetic data but not in the original, $d_{.q}$ and $d_{tq}$ are all zero.
+Similarly when a member of $q$ is in the SD but not in the original, $d_{.q}$ and $d_{tq}$ are all zero.
 The two tables can be written as shown in Table 1, where the total records in the original data is $N_{d}$, made up of $N_{d~only}$ and $N_{d~both}$.
-The equivalent totals for the synthetic data are $N_{s}$, $N_{s~{}only}$ and $N_{s~{}both}$.
+The equivalent totals for the SD are $N_{s}$, $N_{s~only}$ and $N_{s~both}$.
 only in original in both only in synthetic Total
@@
 Totals 0 $N_{s~both}$ $N_{s\>only}$ $N_{s}$
-Table 2: Notation for tables from quasi-identifier ($q$) and target ($t$) from original (upper table) and synthetic data (lower table).
-To calculate the % of records in the original and synthetic data we need:
-%$\%~{}Unique~{}in~{}Original=UiO=100\sum{(d_{.q}|d_{.q}=1})/N_{d}.$ (1)
-%$\%~{}Unique~{}in~{}Synthetic=UiS=100\sum{(s_{.q}|d_{.q}=1})/N_{d}.$ (2)
-The intruder has information about the keys for an individual in the real data that they attempt to identify in the synthetic data.
-They first attempt to find them in the synthetic data, abd the % found is:
-%$\%~{}Unique~{}in~{}Original~{}in~{}Synthetic=UiOiS=100\sum{(d_{.q}=1|s_{.q}=1\land d_{.q}>0})/N_{d}.$ (3)
-Some of these records would not be unique in the synthetic data, restricting to such records gives:
-%$\%~{}replicated~{}Uniques=repU=100\sum{(s_{.q}|d_{.q}=1\land s_{.q}=1)}/N_{d}.$ (4)
+Table 3: Notation for tables from quasi-identifier ($q$) and target ($t$) from original (upper table) and SD (lower table).
+To calculate the % of records in the original and SD we need:
+$\%~Unique~in~Original=UiO=100\sum{(d_{.q}|d_{.q}=1)}/N_{d}.$ (1)
+$\%~Unique~in~Synthetic=UiS=100\sum{(s_{.q}|d_{.q}=1)}/N_{d}.$ (2)
+The intruder has information about the keys for an individual in the real data that they attempt to identify in the SD. They first attempt to find them in the SD, and the % found is:
+$\%~Unique~in~Original~in~Synthetic=UiOiS=100\sum{(d_{.q}=1|s_{.q}=1\land d_{.q}>0)}/N_{d}.$ (3)
+Some of these records would not be unique in the SD, restricting to such records gives:
+$\%~replicated~Uniques=repU=100\sum{(s_{.q}|d_{.q}=1\land s_{.q}=1)}/N_{d}.$ (4)
 To find an attribute from a set of keys, it is necessary to examine the distribution of $s_{tq}$ for groups defined by $q$.
-We define column proportions for the original and synthetic data as $pd_{tq}=d_{tq}/d_{.q}$ and for the synthetic as $ps_{tq}=s_{tq}/s_{.q}$.
+We define column proportions for the original and SD as $pd_{tq}=d_{tq}/d_{.q}$ and for the synthetic as $ps_{tq}=s_{tq}/s_{.q}$, and noting that we set $ps_{tq}$ and $pd_{tq}$ to zero whenever their denominators, $d_{.q}$ and $s_{.q}$ are zero.
 Returning to the scenario described in Section 3.1, we must first define a measure of attribute disclosure for the original data.
 This is based on the concept of l-diversity [12] that requires that each set of records defined by $q$ has at least $l(\geq{2}$ distinct values of the target.
-A data set is l2-diverse for $q$ and $t$ if all records for every $q$ have the same level of $t$ [^11] This could be generalised to $l>2$, but in practice the levels of targets are not generally exchangeable and a more practical approach would be to aggregate levels for certain targets.
+A data set is l2-diverse for $q$ and $t$ if all records for every $q$ have the same level of tt [^17] This could be generalised to $l>2$, but in practice the levels of targets are not generally exchangeable and a more practical approach would be to aggregate levels for certain targets.
 An attribute disclosure measure for the original data can be defined as % Disclosive in Original :
 $Dorig=100\sum^{q}{\sum^{t}{(d_{tq}|pd_{tq}=1)}}/N_{d}.$ (5)
-The equivalent measure for the synthetic data, taken as if it were real, becomes :
+The equivalent measure for the SD, taken as if it were real, becomes :
 $Dsyn=100\sum^{q}{\sum^{t}{(s_{tq}|ps_{tq}=1)}}/N_{s}.$ (6)
-An intruder with access only to the synthetic data, but with knowledge of $q$ from one or more individuals in the original, would look them up in the synthetic data.
-Some of their $q$ levels be key combinations that do not appear in the synthetic data leaving the proportion that do appear as $iSO$ (in in Synthetic Original)
+An intruder with access only to the SD, but with knowledge of $q$ from one or more individuals in the original, would look them up in the SD.
+Some of their $q$ levels be key combinations that do not appear in the SD leaving the proportion that do appear as $iSO$ (in in Synthetic Original)
 $iSO=100\sum^{q}{\sum^{t}{(d_{tq}|s_{tq}>0)}}/N_{d}.$ (7)
-A level of $q$ from an original record may identify more than one target in the synthetic data, or identify the wrong target.
-To exclude these we require the records to be Disclosive in the Synthetic data and Correct when checked with the original giving:
+A level of $q$ from an original record may identify more than one target in the SD, or identify the wrong target.
+To exclude these we require the records to be Disclosive in the SD and Correct when checked with the original giving:
 $DiSCO=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/N_{d}.$ (8)
@@
 As we comment above the intruder would not be able to tell if records were identified as $DiSDiO$ rather than $DiSCO$, so we prefer the latter measure.
-However, the intruder can identify when the apparently disclosive record is not unique in the synthetic data.
-This restriction can be imposed by requiring the denominator in the synthetic data to exceed a 1, as described and discussed in Section 5.
+However, the intruder can identify when the apparently disclosive record is not unique in the SD.
+This restriction can be imposed by requiring that the denominator for disclosive records in the original does not exceed a 1, as described and discussed in Section 5.
 Appendix 2: CAP measures
-The measures $baseCAPd$, $DCAP$, $TCAP$ are calculated by the function disclosure and are stored in the component allCAPs of the output object of class disclosure.
+The measures $baseCAPd$, DCAPDCAP, TCAPTCAP are calculated by the function disclosure and are stored in the component allCAPs of the output object of class disclosure.
 This is printed when the parameter to.print includes allCAPs.
@@
 Averaging this over all observations gives
-$baseCAPd=\sum{(pd_{t.})^{2}}$
+$baseCAPd=\sum{(pd_{t.})^{2}}.$
 The % of $t$ that would be correctly predicted from $q$ for someone with access to the original data is $CAPd$:
 $CAPd=\sum\limits_{tq}{(pd_{tq}d_{tq})}/N_{d},$
-and the equivalent measure for the synthetic data treated as if it were the original is
-$CAPd=\sum\limits_{tq}{(ps_{tq}}s_{tq})/N_{s}.$
-The measure DCAP is the percentage of $t$ correctly predicted in the synthetic data $q$. This gives
-$DCAP=\sum\limits_{tq}{(ps_{tq}}d_{tq})/N_{d}$
-The $DCAP$ score can be expressed by scaling it from $baseCAPD$ to 1[10, 14], although this can result in some negative values.
-The $TCAP$ measure as defined by [10] uses the same count of disclosive records as $DISCO$, but uses as denominator the number of records that have keys represented in both the synthetic and the original data.
-It can also be scaled from $baseCAPD$ to 1, and as [10] demonstrate this measure can also take negative values.
-R> print(t5, to.print = "allCAPs")
-Disclosure measures from synthesis for 5000 records in original data.
-CAP measures for the target depress with keys
-sex age region placesize
-  baseCAPd  CAPd  CAPs  DCAP  TCAP
-1     9.81 74.15 69.78 16.39 14.70
-2     9.81 74.15 69.36 17.45 16.03
-3     9.81 74.15 69.24 16.20 14.21
-4     9.81 74.15 69.87 15.92 14.40
-5     9.81 74.15 69.00 16.17 14.72
-As expected the lower denominator for $TCAP$ gives values higher than the equivalent $DiSCO$ values for the same example given in Section 3.3.
-Appendix 3:: Code for analysis of the Adult data set.
-library(arules)
-data(AdultUCI)
-names(AdultUCI) <- gsub("-",".", names(AdultUCI))
-dim(AdultUCI)
-##------ drop one of two identical education variables---------------------------
-adultsample <- AdultUCI[,-c(5)]
-names(adultsample)[-c(1,6,8,9)]
-names(synadult\$syn)
-##------ native.country moved to end as many levels that slow down synthesis-----
-synadult <- syn(adultsample, visit.sequence = c(1:12,14,13) ,seed = 7868)
-##------ get basic disclosure and make Figure 2----------------------------------
-dd1 <- multi.disclosure(synadult,adultsample,
-   keys = c("age" , "occupation","race", "sex" ))
-pdf("vignettes/fig2dis.pdf",height=300, width=500)
-print(dd1)
-dev.off()
-##------ exclude some target levels and missing values for targets---------------
-dd2 <- multi.disclosure(synadult,adultsample,
-    print.flag = FALSE, not.targetslev = c(rep("",5),"0","0","","United-States",""),
-    keys = c("age" , "occupation","race", "sex" ))
-##------ exclude some target levels and missing values for targets---------------
-dd3 <- multi.disclosure(synadult,adultsample, print.flag = FALSE,
-   not.targetslev = c(rep("",5),"0","0","","United-States",""),
-    keys = c("age" , "occupation","race", "sex" ))
-    dd1 <- multi.disclosure(synadult,adultsample, keys = c("age" , "occupation","race", "sex" ))
-##-------restrict to denominators of 1 -----------------------------------------
-dd4 <- multi.disclosure(synadult,adultsample, denom_lim =1,
-    not.targetslev = c(rep("",5),"0","0","","United-States",""),
-    usekeysNA = FALSE, usetargetsNA = FALSE,
-    exclude_ov_denom_lim = TRUE,
-    keys = c("age" , "occupation","race", "sex" ))
-##----------just denominators not others---------------------------------------
-dd5 <- multi.disclosure(synadult,adultsample, denom_lim =1,
-    exclude_ov_denom_lim = TRUE,
-    keys = c("age" , "occupation","race", "sex" ))
-##-----------make table 1------------------------------------------------------
-library(stringr)
-cols12 <- dd1\$attrib.table[,1:2]
-  dimnames(cols12)[[1]] <- word(dimnames(cols12)[[1]],2,2)
-cols34 <- dd2\$attrib.table[,1:2]
-  dimnames(cols34)[[1]] <- word(dimnames(cols34)[[1]],2,2)
-cols56 <- dd3\$attrib.table[,1:2]
-  dimnames(cols56)[[1]] <- word(dimnames(cols56)[[1]],2,2)
-cols78 <- dd4\$attrib.table[,1:2]
-  dimnames(cols78)[[1]] <- word(dimnames(cols78)[[1]],2,2)
-cols910 <- dd5\$attrib.table[,1:2]
-  dimnames(cols910)[[1]] <- word(dimnames(cols910)[[1]],2,2)
-table.adult <- data.frame(cols12[order(dimnames(cols12)[[1]]),],
-    cols34[order(dimnames(cols34)[[1]]),],
-    cols56[order(dimnames(cols56)[[1]]),],
-    cols78[order(dimnames(cols78)[[1]]),],
-    cols910[order(dimnames(cols910)[[1]]),])
-table.adult
-##--------------------- check 2way for one target------------------------------
-dis1 <- disclosure(synadult,adultsample, target = "workclass",
-    keys = c("age" , "occupation","race", "sex" ))
-print(dis1, to.print = c("attrib","check\_2way"))
+and the equivalent measure for the SD treated as if it were the original is
+$CAPs=\sum\limits_{tq}{(ps_{tq}s_{tq})}/N_{s}.$
+The measures $DCAP$ and $DCAP_{alt}$ are each based on the percentage of synthetic records where $t$ is correctly predicted in the SD This gives
+$DCAP=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{sboth}~~~DCAP_{alt}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{s}~~~DCAP_{alt}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{d}$
+The $TCAP$ and $TCAP_{alt}$ measure use the same count of disclosive records as $DiSCO$ in their numerators, but use different denominators.
+$TCAP_{alt}$ was mistakenly quoted in earlier versions of this vignette as being the measure proposed in [11], using $N_{s~both}$ as its denominator:
+$TCAP_{alt}=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/N_{s~both}.$ (10)
+but the measure used in [11, 10] had a smaller denominator restricted to all counts in the original datam with keys corresponding to those predicted with certainty from the SD, for any target:
+$TCAP=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/\sum^{q}(d_{.q}\sum^{t}{(ps_{tq}=1)}),$ (11)
+where the summation over $t$ in the denominator indicates columns of the $tq$ tables where any target is predicted with certainty from the SD,
+See section 3.3.2 for examples of these measures.

### v5

@@
 Abstract
-This paper explains how synthpop package 111see www.synthpop.org.uk  [^2] The new version 1.9-3 is currently available on Github at https://github.com/gillian-raab/synthpop has been extended to include functions to calculate measures of identity and attribute disclosure risk for synthetic data that measure risks for the records used to create the synthetic data.
-This paper is also appears as a vignette as pat of the versions of the package from 9.3 onwards.
+This paper explains how synthpop package 111see www.synthpop.org.uk has been extended to include functions to calculate measures of identity and attribute disclosure risk for synthetic data that measure risks for the records used to create the synthetic data.
 The basic function disclosure calculates identity disclosure for a set of quasi-identifiers (keys) and attribute disclosure for one variable specified as a target from the same set of keys.
@@
 We recommend two measures: $RepU$ (replicated uniques) for identity disclosure and $DiSCO$ (Disclosive in Synthetic Correct Original) for attribute disclosure.
-Both are expressed a % of the original records and each can be compared to similar measures calculated from the original data.
+Both are expressed a % of the number of original records and each can be compared to similar measures calculated from the original data on its own.
 Experience with using the functions on real data found that some apparent disclosures could be identified as coming from relationships in the data that would be expected to be known to anyone familiar with its features.
 We flag cases when this seems to have occurred and provide means of excluding them.
+This paper was written as a vignette for the R package synthpop version 1.9-3.
 1 Introduction
@@
 a large data set with only a few categorical variables, the disclosure risk from the original may already be low.
-To evaluate the protection from disclosure afforded by a synthesis method, the risks for SD must be compared with equivalent risks for the GT.
+To evaluate the protection from disclosure afforded by a synthesis method, the risks for SD must be compared with equivalent risks for the original ground truth data (GT).
 We consider two types of disclosure risk:
@@
 This is in contrast to incomplete or partial synthesis where only some variables are replaced.
-data because there is no one-to-one correspondence between records in the original and SD.
+data because there is no one-to-one correspondence between records in the GT and the SD.
 But it may still be of interest since it is an important factor for attribute disclosure.
-  • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the original data.
+  • attribute disclosure:This refers to the ability to find out from the keys something, not previously known, for an attribute associated with a record in the GT.
 The measures we describe here are appropriate for fully SD where all items in all records are replaced by synthetic values.
@@
 The disclosure risk posed by SD can be reduced by using techniques from statistical disclosure control (sdc), such as aggregation of categories, smoothing of numeric values or removal of replicated uniques.
-These methods can be used to reduce disclosure risk by modifying the original data before synthesis, or the SD before its is released.
+These methods can be used to reduce disclosure risk by modifying the GT before synthesis, or the SD before its is released.
 Some such methods are already available in the synthpop package.
@@
 We will not discuss these methods here, but note that we could use the metrics proposed here to evaluate disclosure risks for DP SD.
+The Appendix introduces the notation we will use to define the disclosure measures that can be computed by the disclosure functions in the synthpopsynthpop package.
+Our approach involves creating a composite variable, $q$, from the quasi-identifiers (keys) for the original and synthetic data.
+Identity disclosure measures are calculated from univariate tables of $q$ for the GT and the SD.
+Attribute disclosure measures for a target variable $t$ are obtained by tabulating $t$ against $q$ for the GT and the SD.
+For our initial example in Section 2 and most of the later development we assume thet the number of records created in the SD ($N_{s}$) is the same as the number in the GT ($N_{d}$).
+The exception is Section 3.3.2 where we show how different proposed measures for attribute disclosure are affected by reducing the number of synthetic records ($N_{s}$) that are released.
 2 A simple example
@@
 If the parameter targets is not specified, all the variables in the SD that are not part of keys are used as targets.
-The identity disclosure measures are UiO for original and repU for synthetic, and for attribute disclosure Dorig for original and DiSCO for synthetic.
+The identity disclosure measures are UiO for the original and repU for synthetic, and for attribute disclosure Dorig for original and DiSCO for synthetic.
 These and other measures will be explained in Sections 3.2 and 3.3.
@@
 Note that the variable income has values -8 that indicate not applicable, and the synthesis allows for this.
-The synthetic data object s1555an object of class synds has a component syn that is a single synthetic data set.
+The synthetic data object s1444an object of class synds has a component syn that is a single synthetic data set.
 The disclosure functions can also be used with synthetic data created by other methods either as single synthetic data sets or lists of repeated syntheses from the same original.
@@
 In this case 18 groups are formed from the numeric values, plus two groups from the -8 code and the missing values, giving a total of 20 We first use multi.disclosure to create a multi.disclosure object for these keys, and we print out its identity component.
-$\begin{Schunk}$
-$\begin{Sinput}$
 R> library("synthpop")
@@
 R> print(t1,to.print = "ident")
-[…]
-$\begin{Soutput}$
 Disclosure risk for 5000 records in the original data
@@
 For synthetic ( repU ) 14.86 %.
-[…]
-[…]
 The measure $UiO$ (Unique in Original) shows that 48% of the original records would have unique combinations of these 4 keys.
@@
 Here we will introduce the measures by an example.
-Formal definitions with notation and formulae are in Appendix 1.
+Formal definitions with notation and formulae are in the Appendix.
 The first step in evaluating disclosure risk, as described here, is to identify a set of keys that might be expected to be known to an intruder.
@@
 Based on this idea, the percentage of records for which the keys identify just one individual give identity disclosure measures.
-Tables of $q$ values are produced from the synthetic and the original data.
+Tables of $q$ values are produced from the SD and the GT.
 $UiO$ and $UiS$ are the percentages of records with keys where the table count is 1.
@@
 Some records will not be in the SD and $UiOiS$ (Unique in Original in Synthetic) gives the % that would be found.
-These records are then checked for uniqueness in the synthetic data giving, $repU$ is the percentage of unique original records that are also unique in the SD.
+These records are then checked for uniqueness in the synthetic data giving, $repU$ is the percentage of unique records in GT that are also unique in the SD.
 The percentage $repU$ has been used as a disclosure measure to evaluate SD by [9] and by [17]  [^8] Jackson et al.
 in [9] argue that the denominator for $repU$ should be $N_{s}$ rather than $N_{d}$.
-This is inappropriate because our scenario is to consider the risk to the original data..
+This is inappropriate because our scenario is to consider the risk to the original data, as we discuss for attribute disclosure in Section 3.2.2..
 Replicated uniques are used in synthpop as part of the statistical disclosure control function, sdc, that includes the option of reducing disclosure risk by removing them from the SD.
@@
 Each are steps towards calculating $repU$.
-$\begin{Schunk}$
-$\begin{Sinput}$
 R> s5 <- syn(ods, seed = 8564, m = 5, print.flag = FALSE)
@@
 R> print(t5, to.print = c("ident"))
-[…]
-$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
@@
 5 48.38 36.14 22.00 14.62
-[…]
-[…]
 3.3 Attribute disclosure measures
-The attribute disclosure measure for the SD that is displayed in Figure 1 is our preferred measure, $DiSCO$ [16].
-In section 3.3.1 we explain it via the example above.
-Other related disclosure measures, $DCAP$ and $TCAP$, have been proposed [6, 23, 22, 11, 10].
-The disclosure function also computes these measures.
-Section 3.3.2 describes them and discusses circumstances where they differ from $DiSCO$.
-The reason for our preference for $DiSCO$ compared to other proposed methods is that it is a mesure computed for all the records in the original data.
-The other measures thus do not include aspects of disclosure, for example that some records in the original not appearing in the SD, that are valid contributions to the disclosure reduction effected by the synthesis.
-The measure we describe as $DCAP$ is readily adapted to apply to all original records by replacing $N_{s}$ in its denominator with $N_{d}$.
-When the synthesis creates SD of the same size as the original $DCAP$ and $Dorig$ are identical.
-But, as we show in Figure 2, $DCAPorig$ assesses disclosure appropriately when synthetic data with fewer records than the original are created.
-This measure is added to the component attribattrib of the disclosure object created by the disclosure function.
-The two measures $DiSCO$ and $DCAPorig$ can be considered as the extremities of a range of a more general disclosure measure when a record is considered disclosive if its probability of correct attribution reaches a threshold tautau.
-DiSCODiSCO corresponds to a tautau of 1.0 while $DCAPorig$ corresponds to a tautau of zero.
+The attribute disclosure measures displayed in Figure 1 are our preferred measures, $DiSCO$, for the SD and $Dorig$ for the GT [16].
+In section 3.3.1 we explain them via the example above.
+Other related disclosure measures, $DCAP$ and $TCAP$ and their variants, have been proposed .
+We prefer $DiSCO$ over $TCAP$ and its variants [11, 10] because it measures all the ways synthesis is effective in reducing attribute disclosure.
+A further measure $DCAP_{d}$, a modification of [6, 22], also measures all these aspects.
+The measures $DiSCO$ and $DCAP_{d}$ are included in the attribattrib component of the output from the disclosure function.
+For completeness, all the other measures are computed by disclosuredisclosure and returned in the $allCAPs$ component of the disclosure object.
+The two measures $DiSCO$ and $DCAP_{d}$ are the only two measures that show how the disclosure risk is reduced by releasing SD with fer records than GT, see Figure2.
+They can be considered as the extremities of a range of a more general disclosure measure when a record is considered disclosive if its probability of correct attribution reaches a threshold tautau.
+DiSCODiSCO corresponds to a tautau of 1.0 while $DCAP_{d}$ corresponds to a tautau of zero.
 This more general approach may be included in future releases of synthpopsynthpop.
-Appendix 1 gives formal definitions of all disclosure measures discussed here.
-3.3.1 Our preferred measure: Disclosive in Synthetic Correct in Original ($DiSCO$)
+Section 3.3.2 discusses all these other measures and formal definitions of all measures are in the Appendix.
+3.3.1 Disclosive in Synthetic Correct in Original: DiSCODiSCO
 We approach disclosure risk from the point of view of an intruder with access to the SD and to certain attributes (quasi-identifiers) known for one or more individuals in the original data.
@@
 Modelling what an intruder might do we calculate the following measures, each of which is a proportion of the original records:
-  • Look up $q$ for each original record in the SD. The proportion found becomes $iS$ (in Synthetic).
-  • Check if all records with the same $q$ have the same level of the target $t$. The proportion passing this further test becomes $DiS$ (Disclosive in Synthetic).
+  • Look up $q$ in the SD for each original record in the GT. The proportion found becomes $iS$ (in Synthetic).
+  • Check if all records with the same $q$ have the same level of the target $t$. The proportion of all GT records passing this further test becomes $DiS$ (Disclosive in Synthetic).
   • Then check if these apparent disclosures correspond to the value of $t$ in the original data.
@@
 A further measure $DiSDiO$ (Disclosive in Synthetic Disclosive in Original) restricts the score to those disclosive in the original.
-$\begin{Schunk}$
-$\begin{Sinput}$
 R> print(t5, to.print = c("attrib"))
-[…]
-$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
@@
 5  53.3 44.52 63.44 31.46  9.34    16.17   5.80         4       1.23
-[…]
-[…]
 Here we print the table of attribute disclosure measures for the example in the previous section.
@@
 The example in Section 2 is here run restricted to denominators of 1.
-$\begin{Schunk}$
-$\begin{Sinput}$
 R> multi.disclosure(s1, ods, print.flag = FALSE, plot = FALSE,
@@
 +    denom_lim =1, exclude_ov_denom_lim = TRUE)
-[…]
-$\begin{Soutput}$
 Disclosure risk for 5000 records in the original data
@@
 5 workab 1way checks
-[…]
-[…]
 The $DiSCO$ values have decreased, as expected, but $UiO$ and $repU$ have increased and also that $Dorig$ is now the same for all targets and equal to the $UiO$ value before the denominator exclusion.
@@
 3.3.2 Other attribute disclosure measures
-Previous work on attribute disclosure [6, 23] has used the Correct Attribution Probability ($DCAP$) as a disclosure measure for synthetic data.
-This is calculated as the average percentage of records correctly predicted from the synthetic data by a set of keys.
-Two different versions are described to deal with the case when some key combinations in the SD do not occur in the original, so that the attribution probability is undefined.
-In one case, that we will designate as $DCAP$, the undefined attribution probabilities are set to zero and the denominator is the total number of synthetic records.
-The second case $DCAP_{alt}$ [^11] This approach is sometimes referred to as $TCAP_{undefined}$ [23].
-uses the number of synthetic records present in the synthetic data as the denominator, so that $DCAP_{alt}\geq DCAP$.
-The related quantity $TCAP$ restricts the probabilities to those over a certain limit, and most commonly this limit is $1.0$ making the numerator of the $TCAP$ measures identical to that for $DiSCO$.
-One possibility for a denominator would be to use just the number of synthetic records with key combinations in the original, as for $DCAP_{alt}$ [^12] This measure was what was reported as $TCAP$ in Versions of $Synthpop$ before Version 1.9.3.
-But evaluations using $TCAP$ [11, 10] use an even smaller denominator; the number of synthetic records with keys present in the original and where the target probaility is 1.0.
-This can be described as the proportion correctly predicted for all apparently disclosive records; we refer to this measure as $TCAP$ and note that its smaller denominator implies that $TCAP\geq TCAP_{alt}\geq DiSCO$ when the synthetic data set(s) are the same size as the original.
+Previous work on attribute disclosure [6, 22] has used the Correct Attribution Probability ($CAP$) as a disclosure measure for synthetic data.
+This is the proportion for a record in the SD where its $q$ correctly predicts the attribute level of the target, $t$, in the original.
+Two versions are described, each with the total $CAP$ in the numerator, but different denominators.
+Some combinations of the keys, $q$, found in the SD may not exist in the original, so that their $CAP$ is undefined.
+One version, that we will designate as $DCAP_{b}$, ignores the records with undefined probabilities and uses as a denominator $N_{b}$, the number of synthetic records where the value of $q$ occurs in the original.
+[^10] This approach is sometimes referred to as $DCAP_{undefined}$ [22].
+A second version ($DCAP_{s}$) sets the undefined attribution probabilities to zero for records where $CAP$ is undefined and uses as denominator ($N_{s}$) the total number of synthetic records.
+Note that $DCAP_{b}\geq DCAP_{s}$ because $N_{b}\leq N_{s}$.
+If the total mumber of original records is used as the dsenominator we get the measure $DCAP_{d}$.
+This is the same as $DCAP_{s}$ when the SD has the same number of records as the GT, but will be lower when smaller SD samples are generated.
+The related quantity $TCAP$, and its variants.
+restricts the probabilities to those over a certain limit, and most commonly this limit is $1.0$.
+WIth this choice the numerator of the $TCAP$ measures identical to that for $DiSCO$, the proportion of synthetic records correctly predicted with certainty from the SD.
+The same two choices of denominator as were used for $DCAP$ could be used.
+One choice $TCAP_{s}$ would be to use the denominator, $N_{s}$ and the other $TCAP_{b}$ would be to use the denominator, $N_{b}$.
+Note that when $N_{s}=N_{d}$, the measure $TCAP_{s}$ will be identical to $DiSCO$.
+However the $TCAP$ that is defined in [11, 10] does not use either of these denominators, but uses as a denominator $N_{bp}$ the number of synthetic records with $q$ present in the original and where any level of the target $t$ is predicted with certainty, but not necessarily correct.
+Another way of describing this $TCAP$ is as the proportion of apparently disclosive recordsin the SD that are correct in the original.
+The numerators and denominators used for each of these measures are summarised in Table 1.
+Numerator
+Sum of proportions correct Total number predicted with certainty
+Denominator CAP CAP = 1
+$N_{d}$ $DCAP_{d}$ DiSCODiSCO
+$N_{s}$ $DCAP_{s}$ $TCAP_{s}$
+$N_{b}$ $DCAP_{b}$ $TCAP_{b}$
+$N_{bp}$ TCAPTCAP
+Table 1: Summary of numerators and denominators for $DCAP$ and TCAPTCAP measures.
 All of these measures are shown in Table 2 for 3 of the targets in the synthesis of the 5,000 records from the example above.
@@
 For this case the ordering of the disclosive risk by target is consistent for all measures.
-For the three measures that share the same numerator ($DiSCO$, $TCAP_{alt}$ and $DCAP$) we see that the reduced denominators have a large effect in this example, so that $TCAP$ is around three times $DiSCO$.
-DiSCO DCAP DCAPalt TCAPalt TCAP
-ls 12.63 21.97 29.57 17.00 36.71
-depress 9.21 16.23 21.85 12.40 28.57
-income 6.16 11.34 15.27 8.29 19.96
-Table 1: Different disclosure measures (%) for 3 targets from the example above when number of synthetic records equals the number of original records (5000).
+Because the number of records in the GT and the SD are the same, $DCAP_{s}$ is identical to $DCAP_{d}$ and $TCAP_{s}$ is identical to $DiSCO$.
+For measures that share the same numerator, Table 1, we see that the reduced denominators have a large effect in this example, so that $TCAP$ is around three times $DiSCO$.
+$DCAP_{s}$ and $DCAP_{d}$ $DCAP_{b}$ $TCAP_{d}$ and DiSCODiSCO $_{b} TCAPTCAP$
+ls 21.97 29.57 12.63 17.00 36.71
+depress 16.23 21.85 9.21 12.40 28.57
+income 11.34 15.27 6.16 8.29 19.96
+Table 2: DCAP and TCAP variants (%) for 3 targets from the example above when number of synthetic records equals the number of original records (5000). Average results for 5 SDs.
+For all the examples we have examined so far SD has been generated with sample size $N_{s}$ the same as the number of original records $N_{d}$; $N_{s}=N_{d}$.
 Another way that the disclosure risk to the original records can be reduced is to release synthetic data sets that are smaller than the original.
 This is the approach that has been used when UK Census data has been released [2] or for low-fidelity Census data from Scotland [^11] https://www.researchdata.scot/news-and-insights/synthetic-census-data-now-available-for-research/ accessed 7/5/25.
+Figure 2: Plot of disclosure measures for 3 targets for synthetic data of size ($k$) from 250 to 5000.
+Note that the legends apply to all figures, with type of line indicating the numerators and colours the denominators of each measure.
 Figure 2 illustrates how the different disclosure measures assess the risks in this example for sizes of SD from 5000, as the original, to only 250 records.
-Of the measures in Table 1, only $DiSCO$ and $DCAPorig$ adjust the attribute disclosure risk for the reduced number of records in the SD.
+Of the measures in Table 1, only $DiSCO$ and $DCAP_{d}$ adjust the attribute disclosure risk correctly for the reduced number of records in the SD.
 The $TCAP$ measure is almost independent of the size of the SD.
-The other measures all give misleadingly high disclosure risks for small samples.
-Figure 2: Plot of disclosure measures for synthetic data of size ($k$) from 250 to 5000.
-Figure 2 displays
+The other measures all give misleadingly high disclosure risks for small samples.We can see how the pairs of measures ($DiSCO$, $TCAP_{s}$) and ($DCAP_{d}$$DCAP_{s}$) diverge for $N_{s}$ values lower than $N_{d}$.
+Referring back to section 3.3.1 we can see how these measures relate to the three bullet points giving the stages by which an intruder takes to identify an attribute disclosure.
+The measure $TCAP$ assesses the intruder’s success rate only at the final bullet point, starting with records in $DiS$.
+$TCAP_{s}$ measures the success of the last two steps, from $iS$, but only $DiSCO$ includes the disclosure risk from all 3 steps.
 4 Identifying disclosure from 1-way and 2-way relationships
@@
 These two aspects are flagged by the values of check_1way and check_2way that are returned as part of objects of Returning to the example in Section 2, we now use the function disclosure to get details for the two targets that were flagged as requiring checking; see Figure 1.
-$\begin{Schunk}$
-$\begin{Sinput}$
 R> d1_workab <- disclosure(s1, ods, print.flag = FALSE, target = "workab",
@@
 R> print(d1_workab, to.print = c("check_1way"))
-[…]
-$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
@@
 exclude.keys and exclude.keylevs
-[…]
-[…]
 We can see that it was the category ”NO” of workab that contributed most to the disclosure risk; most survey respondents (89%) had never worked abroad.
@@
 The $tq$ pairs identified for workab both included the ”NO” level of the target.
-$\begin{Schunk}$
-$\begin{Sinput}$
 R> d1_marital <- disclosure(s1, ods, print.flag = FALSE, target = "marital",
@@
 R> print(d1_marital, to.print = c("check_2way"))
-[…]
-$\begin{Soutput}$
 Disclosure measures from synthesis for 5000 records in original data.
@@
 8       SINGLE|22      5 age               85        91             93.41
-[…]
-[…]
 For the target marital two $tq$ pairs are identified as contributing large denominators.
@@
 The summary function multi.disclosure does not allow target and key specific pairs to be excluded.
-To investigate this it is necessary to examine the output of disclosure151515see code in Appendix 3 for how to do this..
+To investigate this it is necessary to examine the output of disclosure.
 This showed that the largest contribution to check_2way for workclass was due to this being missing when occupation was missing, although other relationships between these two variables also contributed.
@@
 7 Acknowledgement
-Research Data Scotland (https://www.researchdata.scot/) funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package161616from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control.
+Research Data Scotland (https://www.researchdata.scot/) funded Gillian Raab’s time to carry out the research reported here and to expand the capabilities of the synthpop package131313from version 1.8-0 available on CRAN at https://CRAN.R-project.org/package=synthpop to include measures of disclosure control.
 We also thank the Scottish Centre for Administrative Data Research for continue to support the development of the synthpop package since its creation was supported by the ESRC funded SYLLS poject in 2012-14.
 We would also like to thank Timon Huijser, for pointing out errros in a previous version of this vignette.
-Appendix 1: Notation and formal definitions
+Appendix
+Notation and formal definitions
 Before defining the measures of identity and disclosure risk we need to introduce the notation that will be used to calculate them.
@@
 The two tables can be written as shown in Table 1, where the total records in the original data is $N_{d}$, made up of $N_{d~only}$ and $N_{d~both}$.
-The equivalent totals for the SD are $N_{s}$, $N_{s~only}$ and $N_{s~both}$.
+The equivalent totals for the SD are $N_{s}$, $N_{s~only}$ and $N_{b}$.
 only in original in both only in synthetic Total
@@
 Column sums $d_{.q}$ $d_{.q}$ 0 $N_{d}$
-Totals $N_{d\>only}$ $N_{d~both}$ 0 $N_{d}$
+Totals $N_{d\>only}$ $N_{b}$ 0 $N_{d}$
 only in original in both only in synthetic Total
@@
 Column sums 0 $s_{.q}$ $s_{.q}$ $N_{s}$
-Totals 0 $N_{s~both}$ $N_{s\>only}$ $N_{s}$
+Totals 0 $N_{b}$ $N_{s\>only}$ $N_{s}$
 Table 4: Notation for tables from quasi-identifier ($q$) and target ($t$) from original (upper table) and SD (lower table).
@@
 This restriction can be imposed by requiring that the denominator for disclosive records in the original does not exceed a 1, as described and discussed in Section 5.
-Appendix 2: CAP measures
+DCAP and TCAP measures
 The measures $baseCAPd$, DCAPDCAP, TCAPTCAP are calculated by the function disclosure and are stored in the component allCAPs of the output object of class disclosure.
@@
 $CAPs=\sum\limits_{tq}{(ps_{tq}s_{tq})}/N_{s}.$
-The measures $DCAP$ and $DCAP_{alt}$ are each based on the percentage of synthetic records where $t$ is correctly predicted in the SD This gives
-$DCAP=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{sboth}~~~DCAP_{alt}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{s}~~~DCAP_{alt}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{d}$
-The $TCAP$ and $TCAP_{alt}$ measure use the same count of disclosive records as $DiSCO$ in their numerators, but use different denominators.
-$TCAP_{alt}$ was mistakenly quoted in earlier versions of this vignette as being the measure proposed in [11], using $N_{s~both}$ as its denominator:
-$TCAP_{alt}=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/N_{s~both}.$ (10)
-but the measure used in [11, 10] had a smaller denominator restricted to all counts in the original datam with keys corresponding to those predicted with certainty from the SD, for any target:
+The measures $DCAP_{b}$ and $DCAP_{s}$ are each based on the percentage of synthetic records where $t$ is correctly predicted in the SD This gives
+$DCAP_{b}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{b}~~~DCAP_{s}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{s}$
+Another option for the denominator of a $DCAP$ measure is the count of GT records, $N_{d}$, as used for $DiSCO$ in equation (11).
+$DCAP_{d}=\sum\limits_{tq}{(ps_{tq}d_{tq})}/N_{d}$ (10)
+When the SD and GT have the same number of records $DCAP_{s}=DCAP_{d}$. But when the SD has fewer records $DCAP_{d}$ will be reduced.
+The $TCAP$ and $TCAP_{b}$ measures use the same count of disclosive records as $DiSCO$ in their numerators, but use different denominators.
+$TCAP_{b}$ was mistakenly quoted in earlier versions of this vignette as being the measure proposed in [11], using $N_{b}$ as its denominator:
+$TCAP_{b}=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/N_{b}.$ (11)
+but the measure used in [11, 10] had a smaller denominator restricted to all counts in the original data, with keys corresponding to those predicted with certainty from the SD, for any target:
 $TCAP=100\sum^{q}{\sum^{t}{(d_{tq}|ps_{tq}=1)}}/\sum^{q}(d_{.q}\sum^{t}{(ps_{tq}=1)}),$ (12)
-where the summation over $t$ in the denominator indicates columns of the $tq$ tables where any target is predicted with certainty from the SD,
-See section 3.3.2 for examples of these measures.
+where the summation over $t$ in the denominator indicates columns of the $tq$ tables where any target is predicted with certainty from the SD, the quantity $N_{bp}$ defined in section 3.3.2.